### v1

# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs, which are common in domains such as citation networks, social platforms, and knowledge graphs. GNNs are not inherently interpretable and thus, many explanation methods have been proposed. However, existing explanation methods often struggle to generate interpretable, fine-grained rationales, especially when node attributes include rich natural language. In this work, we introduce Logic, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions. Logic projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure. This enables the LLM to reason about GNN internal representations and produce natural language explanations along with concise explanation subgraphs. Our experiments across four real-world TAG datasets demonstrate that Logic achieves a favorable trade-off between fidelity and sparsity, while significantly improving human-centric metrics such as insightfulness. Logic sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context
Peyman Baghershahi1thanks: Equal contribution  Gregoire Fournier [^1] footnotemark: 1  Pranav Nyati2  Sourav Medya1 1University of Illinois Chicago, 2Indian Institute of Technology Kharagpur 1{pbaghe2, gfourn2, medya}@uic.edu; 2pranavnyati26@kgpian.iitkgp.ac.in
1 Introduction
Graph neural networks (GNNs) have witnessed rapid adoption across a wide range of critical real-world applications, including healthcare Zitnik et al. (2018, 2019), drug design Xiong et al. (2021); Liu et al. (2022); Sun et al. (2020), recommender systems Chen et al. (2022), and fraud detection Rao et al. (2021). The high-stakes nature of these domains demands that the predictions made by GNNs be not only accurate but also trustworthy. One powerful approach to building trust in deep learning models is to provide meaningful explanations that justify their predictions Shneiderman (2020). Such explanations, which may highlight important substructures in the input Luo et al. (2020); Schlichtkrull et al. (2021); Yuan et al. (2021), or offer counterfactual scenarios Lucic et al. (2022a); Tan et al. (2022a); Chen et al. (2022). However, generating explanations for GNNs is inherently challenging due to the combinatorial nature of graph data and the joint influence of node attributes and edge connections. These complexities make GNN explainability a non-trivial problem and have led to the development of a wide array of explanation techniques Kakkad et al. (2023a); Guo et al. (2023).
1.1 Need for LLM-based GNN explainers
There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text attribute graphs (TAGs). Essentially, these methods leverage LLMs in different architectural orders to 1) attain rich representations of the input TAGs, task descriptions, and task-specific few-shot examples for the GNN predictors Liu et al. (2024a); Fang et al. (2024), or 2) reason over the TAGs directly Fatemi et al. (2024) or along with GNN-enhanced soft prompts to make predictions Chen et al. (2024a); Tang et al. (2024). However, there is a lack of focus on using LLMs to explain GNNs predictions and clarify the reasoning behind these black-box models.
Traditional GNN explainers aim to identify influential subgraphs or feature subsets that contribute to predictions Ying et al. (2019a); Luo et al. (2020). While effective for graphs with simple node features, these methods struggle in TAGs, where node information is embedded in natural language. Moreover, often these explanations are in the form of subgraphs that are not human-tangible for domain users. LLMs, on the other hand, excel at reasoning over textual content and can generate coherent, human-interpretable rationales. Thus, their integration into GNNs explanation pipelines can improve the interpretability of the GNN models outputs on TAGs.
1.2 Limitations of Existing LLM-based GNN Explainers
The existing frameworks (please see Sec. 2) that use LLMs to build GNN explanations have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainers’ output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or extensive training/fine-tuning of the GNN explainer. (2) Missed GNN internals: Existing approaches fail to directly leverage the rich internal representations of the GNN, making the explanations generic or unfaithful to the internal working of the GNNs. (3) Bias from suboptimal explainers: Involving a pre-existing GNN explainer can bias the reasoning and judgment of the LLM. This is critical when the GNN explainers are suboptimal and they signal noisy information to LLM, while the LLM can potentially better infer and reason the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
1.3 Our Contributions
In this work, we present Logic (Language mOdel-guided GNN Interpretation with Context), a lightweight, post-hoc explanation framework for text-attributed graphs (TAGs) that seamlessly integrates the representational power of GNNs with the reasoning capabilities of LLMs. Our contributions are as follows:
  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by directly using LLMs as interpreters of GNN behavior. Logic generates fine-grained, natural language explanations without relying on handcrafted templates or perturbation-based saliency masks, reducing bias from external explainers.
  • Soft Prompt Integration: Logic aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector. This projection enables GNN features to be treated as soft prompts in a hybrid prompt that blends structural and textual cues, allowing the LLM to reason over the graph’s latent space.
  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Logic produces interpretable subgraphs based on node-level binary support decisions derived from natural language explanations.
  • Training-Agnostic and Plug-and-Play Deployment: Logic does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining. This makes it adaptable for real-world applications.
2 Related Work
We describe the related studies on the integrated LLM and GNN explainer methods. The explanation capability of LLMs has been applied to graph problems in task-oriented contexts such as explainable recommendation systems Li et al. (2025); Ma et al. (2024). There has been attempts to guide GNNs using the explanations generated by LLMs, for example, in He et al. (2024a), an LLM initially generates node predictions and textual explanations on the reasoning behind the predictions. These explanations then guide a second LLM to produce new node features, which augment the original text features as input for a GNN predictor.
LLM explanation to enhance explainers. LLMs have been utilized to enhance the explanations of pre-existing GNN explainers, for example Zhang et al. (2024) uses LLMs to evaluate explanations of an explainer, its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias. Similarly, to improve the understandability of explanations, Pan et al. (2024) first trains a pseudo-label generator LLM that takes explanations from a pre-existing GNN explainers and generates pseudo-labels for them. The generator is tuned in an expert iteration procedure with tailored objectives for faithfulness and brevity, and the optimized generator generates pseudo-labels to fine-tune an explainer LLM to build end-to-end models.
LLM explanation to enhance understanding. One advantage of LLMs is that they can generate natural language explanations of GNNs, which are more interpretable for humans. He et al. (2024b) also proposes a method for generating counterfactual graph explanations for GNNs, which uses a counterfactual autoencoder (CA)—composed of a contrastively pretrained text encoder based on LLM-generated text pairs and a graph decoder—to construct counterfactual graph topologies by LLM-generated counterfactual text pairs. To address hallucination, the framework incorporates a dynamic feedback mechanism to refine the generated counterfactual text pairs by prompting the LLM back again first-round generated counterfactuals. Concurrent to our work Cedro and Martens (2025) suggests a narrative (story-like) explanation for a GNN’s predictions. Given an LLM, the dataset description, and a GNN model’s predictions, along with the explanation generated by a GNN explainer in the form of subgraphs and feature importance matrices, they ask the LLM to narrate the relationship between the neighboring nodes, the edges connecting them, and the node’s features but fall short of having the LLM as a method to generate the explanations. Our work aims at addressing these limitations and producing a GNN internal representation-aware LLM explanation. We present additional related work in Appendix A.
3 Problem formulation
We define a graph as $G=(V,E,X)$, where $V$ is the set of nodes, $E=\{(u,v)\mid u,v\in V\}$ is the set of edges, and $X=\{x_{v}\mid v\in V\}$ is the node feature matrix, with $x_{v}\in\mathbb{R}^{d}$ representing the feature vector of node $v$. We denote by $\mathcal{G}$ the set of graphs in a dataset, and by $\mathcal{V}$ the set of all nodes across the graphs in $\mathcal{G}$.
Definition 1 (Text-Attributed Graph (TAG)).
A Text-Attributed Graph (TAG) is defined as a graph
$G=(V,E,T)$
where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges (undirected or directed), $T:V\rightarrow\mathcal{D}$ is a function that assigns a textual document or sequence to each node, where $\mathcal{D}$ is a corpus of natural language texts (e.g., sentences, paragraphs).
Each node $v\in V$ is thus associated with a text document $T(v)$, and the graph structure $E$ represents semantic, relational, or structural links between nodes. TAGs enable learning over both the textual content and the topology of the graph. In this work, we focus on the node classification task.
Node Classification. Given a dataset $\mathcal{G}$ and a graph $G=(V,E,X)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$.
Our method aims at explaining the node classification predictions of a GNN with local subgraphs, i.e, small subgraphs around the node staged for classification Ying et al. (2019a). More formally:
Definition 2 (Local Factual Explainer).
Given a GNN $\Phi$ and a graph $G=(V,E,T)\in\mathcal{G}$, a local factual explainer $\Psi$ is a mapping from $(\mathcal{V},\mathcal{G})$ to $\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ associated to $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ verifying:
$\Phi(v,\Psi(v,G))=\Phi(v,G)\;\;\forall G\in\mathcal{G},\forall v\in V\cap\mathcal{V}^{\prime}$ (1)
where $\mathcal{S}_{v}(G)\in\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ denotes the set of all subgraphs of $G$ that contain $v$. $\mathcal{V}^{\prime}$ corresponds to the set of nodes over all graphs for which the equation is verified, and in this case, the explanation $\Psi(v,G)$ is called faithful.
Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ verifies Equation 1 for every node. However, such explainers are rarely useful in practice, as they do not isolate concise or interpretable subgraphs. We therefore evaluate explanations based on their size and fidelity, as formally defined in Section 5.1.
Although Definition 2 necessitates faithfulness of the explanation graphs and the existing methods that build explanations for GNNs mostly follow this definition, there are two major problems:
  • It does not necessarily generate concise explanations. Generally, a good factual explanation subgraph is a sparse one Kakkad et al. (2023a); Yuan et al. (2023) which has only the most important features of the graph, i.e., nodes, edges, and node features. This property makes the graph more comprehensible for users.
  • More importantly, this problem definition does not account for human-level understanding of the explanations. For example, retrieving a subgraph as a common form of explanations Ying et al. (2019a); Luo et al. (2020) needs additional information for the domain-experts to understand the results and make them useful in practice.
Nevertheless, these two desirable criteria are indeed compatible and aligned with each other as well as the objective of explainable AI (XAI), which attempts to make black-box AI models more interpretable for humans. Therefore, our goal is to design a method to enhance the understandability of the explanations while making them coherent and concise.
Figure 1: Illustration of Logic’s framework. First, the projector is trained to align GNN node embeddings with the LLM’s embedding space. Next, hybrid prompts are constructed by interleaving projected embeddings (as soft prompts) with natural language tokens. These prompts are then fed into the LLM to produce natural language explanations, which are converted into explanation subgraphs.
4 Our Proposed Method: Logic
To explain the predictions of a trained GNN using an LLM, our method projects node embeddings from the GNN latent space into the LLM embedding space, allowing the LLM to reason about the GNN’s predictions. Rather than relying solely on textual prompts, our approach injects GNN embeddings as soft prompts interleaved with natural language tokens. This enables the LLM to generate context-aware explanations grounded in both the model’s internal representations and human-interpretable features. We present our method in three parts: first, the training of the projector; second, the construction of the hybrid prompt; and third, the generation of the explanation subgraph. An overview of our approach is shown in Figure 1.
4.1 Projection from the GNN
In order for the LLM to explain the GNN model, we aim to build a framework where the LLM is able to understand how the GNN sees each node of the computation tree when classifying a particular node. To that effect, we introduce a projector that maps nodes’ GNN embedding into tokens.
Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer. The projector aims at both aligning its input with the GNN embeddings and grounding its output in the natural language context of the nodes, hence considering two different levels of embeddings.
Generally speaking, this approach has similarities with other multi-modal alignment frameworks. For instance, in Radford et al. (2021), the authors use contrastive training to align image and text embeddings in a shared space, while in Tsimpoukelli et al. (2021), images are projected directly into a frozen language model’s embedding space to enable coherent language understanding. Similarly, our designed projector learns to map GNN embeddings into soft prompt tokens that serve as input to an LLM, effectively bridging the GNN’s latent space and the LLM’s token-level semantics.
Formally, we define a projector as a function:
$\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$
where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings. We write $\Pi(f_{\Phi}(v))=Z_{v}\in\mathbb{R}^{k\times h}$. The projector $\Pi$ is trained to optimize two losses:
  1. 1.
    Context alignment loss: This encourages the average soft token representation to align with the LLM embedding of the natural language text associated with the node:
    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\left\langle\bar{Z}_{v},\;LLM(v)\right\rangle\right]$
    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output (normalized), and $LLM(v)\in\mathbb{R}^{k}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space. The inner product corresponds to cosine similarity.
  2. 2.
    Contrastive loss with GNN embeddings: This encourages soft prompt representations to preserve the similarity structure of the GNN embeddings as follows:
    $\mathcal{L}_{\text{contrast}}=-\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{V}}p_{vu}^{\Phi}\log p_{vu}^{\Pi}$
    where
    $p_{vu}^{\Phi}=\frac{\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(u)\right\rangle/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(w)\right\rangle/\tau\right)}$
    $p_{vu}^{\Pi}=\frac{\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{u}\right\rangle\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{w}\right\rangle\right)}$
    and $\tau>0$ is a temperature hyperparameter. Both similarity scores are computed using cosine similarity between normalized vectors.
    The overall training objective becomes:
    $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$
    where $1\geq\beta\geq 0$ controls the balance between the two losses.
4.2 Prompt Construction
Given a node $u$, the projector outputs a matrix $\Pi(f_{\Phi}(u))=Z_{u}\in\mathbb{R}^{k\times h}$, which we refer to as a soft prompt embedding. This matrix is injected directly into the LLM’s embedding space and interleaved with natural language token embeddings.
For a node $v$ that we wish to explain, we consider its computation tree for GNN $\Phi$, denoted by $\mathcal{T}^{\phi}_{v}$. $\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$, the number of layers of $\Phi$, composed of all paths of length $L$ starting at $v$, with repetitions, concatenated to the root $v$. We construct the LLM input as a sequence of embeddings consisting of:
  • Standard token embeddings for the system prompt and user instruction.
  • A segment labeled “Target Paper Embedding Representation,” where we insert the soft prompt matrix $Z_{v}$ for the target node, bracketed by textual markers.
  • An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is encoded by interleaved natural language markers and its own soft-prompt embeddings.
  • Final instructions text prompting the model to provide a summary and binary support decision for each computation tree node based on the embeddings.
The LLM sees the embedded GNN features as native tokens, enabling reasoning over both GNN’s representation and natural language inputs. We give an example of the resulting prompt in D.
4.3 Explanation Generation
To produce the explanations, the LLM first predicts whether each node in the computation tree supports or does not support the target node’s classification. This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, (iii) Neutral nodes, not assigned to either category. The explanation subgraph is then formed by including all supporting nodes and a controlled number of neutral nodes. The inclusion of hallucinated nodes is proportional to the number of supporting nodes relative to the total size of the computation tree, as described in our size constraint formulation.
For each node $v$, the LLM generates a natural language explanation $E_{v}$ considering the influence of nodes of $\mathcal{T}^{\Phi}_{v}$ with regards to the classification of node $v$. We define a label attribution function based on the parsing of $E_{v}$:
$\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$
such that for each node $u\in\mathcal{T}^{\phi}_{v}$:
  • $\chi(u)=1$ if $u$ is cited in $E_{v}$ as supporting the classification of $v$,
  • $\chi(u)=-1$ if $u$ is cited as not supporting the classification of $v$,
  • $\chi(u)=0$ if $u$ is not mentioned.
We consider the induced partition of $\mathcal{T}^{\phi}_{v}$:
$S^{+}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=1\}$
$S^{-}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=-1\}$
$S^{0}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=0\}$
To address uncertainty and enforce conciseness, we introduce a refinement function $\rho:2^{\mathcal{T}^{\phi}_{v}}\to 2^{\mathcal{T}^{\phi}_{v}}$ such that the final explanation subgraph is:
$S_{v}=\rho(S^{+}_{v},S^{-}_{v},S^{0}_{v})$
[…]
Mitigating Hallucination. LLM-generated explanations may hallucinate, especially when multiple node embeddings are projected within a single prompt. We address this issue through two complementary mechanisms within our framework:
  • Prompt templating with constrained node sets: To reduce hallucinations, we ensure that our prompts contain a fixed set of nodes. We explicitly instruct the LLM to generate explanations only referencing these nodes, limiting the possibility of hallucinating unsupported nodes.
  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes are filtered out in a post-processing step. This filtering process ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
These strategies work together to minimize hallucination and enhance the accuracy and reliability of the generated explanations.
Table 1: Performance comparison with baseline models on the node classification task. Higher fidelity and lower size are better. *The explainer returns all-zero masks on Liar, we consider its output to be the central node only.
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
Node 73.6% 1 77.8% 1 95.0% 1 73.8% 1
Random (50%) 93.5% 18.1 88.9% 19.5 95.0% 19.0 84.4% 7.24
Random (25%) 79.3% 10.0 86.9% 10.0 95.0% 18.13 73.3% 4.03
GNNExplainer 92.0% 18.38 95.4% 9.7 100.0% 12.4 95.5% 5.63
PGExplainer 86.5% 16.2 86.2% 21.2 95.0%* 1.00* 89.0% 5.73
TAGExplainer 60.0% 1.30 80.0% 1.20 95.0%* 1.00* 73.8% 1.52
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.07
Our method 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 1.30
5 Experiments
Our method, Logic, acts as a light-weight framework to generate explanations of Gnns especially for TAGs. We demonstrate the effectiveness of Logic in various settings. Our code is publicly available at: https://github.com/pbaghershahi/LOGIC.git.
5.1 Evaluation metrics
We use the following metrics in the experiments:
1) Fidelity: Fidelity captures how well an explainable model reproduces or approximates the GNN model logic through its prediction.
Mathematically, the fidelity of the explainer $\Psi$ of GNN $\Phi$ on dataset $\mathcal{G}$ over $\mathcal{V}$ nodes is:
$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathds{1}(\Phi(v,\Psi(v,G))=\Phi(v,G))$
2) Explanation size: The explanation size measures the complexity or compactness of the explanation subgraph produced by the explainer $\Psi$. Specifically, it captures the number of nodes in the explanation for a given prediction. Smaller explanation sizes are generally preferred, as they are easier to interpret. The explanation size is:
$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|\Psi(v,G)|$
5.2 Experimental setup
Datasets. We use real-world TAG datasets to evaluate the performance of our method: Cora Sen et al. (2008) - a co-citation network of research papers, WikiCS Mernyei and Cangea (2022) - a network of related Wikipedia articles, the Amazon-Product Feng et al. (2024) dataset - an E-commerce product co-purchase network, and the Liar Wang (2017) dataset - a fake news detection knowledge-graph. Additional details about the processing and relevance of the datasets are provided in Appendix B.
Models. We consider a GCN Kipf and Welling (2017a) as the base GNN model for the node classification task for all datasets. We train a model consisting of three convolutional layers, with the last one being the classification layer. The training/validation/testing split is $60\%/20\%/20\%$ for all the datasets. We give additional information about the GCN training in Appendix C. We use Meta’s Llama-3.1-8B-Instruct from the HuggingFace library as the frozen LLM in our experiments. We run all experiments on NVIDIA-GeForce-RTX-3090 GPUs, 24GB RAM.
Figure 2: A visualization of the explanations on the Amazon dataset. Left shows the explanations subgraph generated by GNNExplainer. Right shows the explanation generated by Logic. The node marked by a bold circle (product #7) is the node of interest for the explanation. Nodes with the same color are from the same category. Although both models have similar quantitative performance (in terms of size and fidelity), the LLM-generated explanations are textually clarified, making them more human-understandable.
Baselines. To compare our method’s performance with existing approaches, we consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph. This represents the minimal possible context and serves as a lower bound on explanation size. (2) Random: We introduce a randomized baseline that selects a subgraph around each target node as its explanation subgraph. Specifically, for each node, we sample a subgraph of half the size of its computation tree. The results are averaged over 5 samples. (3-5) Graph Explainers: We compare Logic with GNNExplainer Ying et al. (2019a), PGExplainer Luo et al. (2020), subgraph-based explainers for node classification tasks, which do not involve any LLM-based component.
We add the recent LLM-based graph explanation method, TAGExplainer Pan et al. (2024). To the best of our knowledge, TAGExplainer is the only existing LLM-based explanation approach specifically designed for TAGs.
5.3 Quantitative Retults
Table 1 presents the results. One major observation is that on all the datasets, the Node baseline (target node as the singleton explanation subgraph) is considerably powerful, denoting how informative the textual attributes of each node are to guide the GNN’s predictions. Assessing based on fidelity, GNNExplainer, achieves the best performance over most of the datasets, except for Cora, on which Logic achieves better results. However, as discussed before, fidelity alone does not reflect the superiority of a method. Essentially, conventional GNNs (GCN in our case) score a sample, i.e., a node, based on the combination of its own information and the aggregated information of its neighbors, meaning that the context surrounding a node highly impacts its prediction score. Therefore, a GNN explainer that explains a target node with its whole induced subgraph achieves 100% fidelity. Meanwhile, the majority of nodes in the same neighborhood of graphs with a high homophily ratio belong to the same category. Thus, selecting a comparably large number of neighbors of a target node on these types of graphs can most probably guarantee the same representation for the target node as if the whole induced subgraph is selected.
Logic achieves a strong balance between fidelity and explanation size across all four datasets. Notably, Logic matches or exceeds the fidelity of traditional explainers like GNNExplainer, outperforming it on Cora, while producing substantially smaller explanation subgraphs, particularly on Liar and Amazon. This result is especially significant for TAGs, where textual node attributes carry rich semantic information; Logic effectively distills these semantics into compact, interpretable explanations without compromising on alignment with the original GNN predictions. In contrast, baseline methods such as Random or Node either inflate the explanation size or overly restrict the context. These results validate our design choice of using LLMs guided by projected GNN embeddings to construct context-aware yet concise explanation subgraphs, offering a superior trade-off between interpretability and predictive fidelity.
We present additional experiments for different GNN architectures and LLMs in Appendix E.
Table 2: Expert qualitative evaluation of the explanations from Logic and GNNexplainers on Cora. $M_{1}$ denotes the mean scores for GNNexplainer and $M_{2}$ the mean scores for Logic.
Dimension $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta(M_{2},M_{1}$)
Understandability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Trustworthiness $2.7_{\scriptstyle\pm 0.8}$ $3.1_{\scriptstyle\pm 0.8}$ $0.4_{\scriptstyle\pm 0.8}$
Insightfulness $3.3_{\scriptstyle\pm 1.1}$ $3.1_{\scriptstyle\pm 1.1}$ $0.2_{\scriptstyle\pm 1.1}$
Satisfaction $2.2_{\scriptstyle\pm 0.9}$ $2.7_{\scriptstyle\pm 0.9}$ $0.5_{\scriptstyle\pm 0.9}$
Confidence $0_{\scriptstyle\pm 0.9} 3.3$$3_{\scriptstyle\pm 0.9}$ $0.3_{\scriptstyle\pm 0.9}$
Convincingness $3.3_{\scriptstyle\pm 1.2}$ $3.7_{\scriptstyle\pm 1.2}$ $0.4_{\scriptstyle\pm 1.2}$
Communicability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Usability $3.3_{\scriptstyle\pm 0.4}$ $3.5_{\scriptstyle\pm 0.4}$ $0.2_{\scriptstyle\pm 0.4}$
5.4 Qualitative Evaluation
Results on Cora. We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix F for a full description of the evaluation dimensions. An expert reviewed the explanations from GNNexplainer and from our method over 10 node classification explanation tasks, and the evaluation framework is developed in Appendix G. The results are presented in Table 5.3. We observe that our method appeared more trustworthy and convincing to our expert, which may be explained by the semantic analysis abilities of the LLM of our graph explaining method. Concurrently, the explanations were also less insightful, as this may have limited the LLM in its exploration of topics and application not directly limited to the subject of the research paper.
Visualization on Amazon dataset. Figure 2 illustrates a representative example of Logic output on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation. In this instance, the target (bold) node corresponds to a product (ID 7), predicted to belong to the “Sports & Outdoors” category. Logic analyzes the local neighborhood of the target node and translates the latent GNN-based representations into natural language explanations by leveraging the LLM’s language generation capabilities, interleaved with soft GNN prompts.
Logic produces node-wise rationales from the computation tree of the target node, explicitly stating for each neighbor a brief semantic summary along with a binary support judgment (YES/NO) regarding its influence on the target’s classification.
Table 3: Performance comparison for the ablation study. LLM denotes prompting the LLM with words only and no projected embeddings. LLM+Pr is LLM with a projector. LLM with projector and post-processing (denoted by LLM+Pr+Po) is our method.
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.2
LLM+Pr 92.2 % 16.0 80.1% 3.7 95.0% 10.1 86.5% 1.7
LLM+Pr+Po 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 4.9
5.5 Ablation Study
To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing. The results are in Table 3. We observe that both fidelity and explanation size drop across all datasets when we replace the words in the prompt with the projected GNN embeddings. This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning. We also observe that adding the post-processing step increases fidelity, at a reduced cost in size compared to the method without a projector, on the WikiCS dataset. From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
6 Conclusions
In this work, we propose Logic, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs. By projecting GNN embeddings into the LLM space, Logic enables context-aware, natural language rationales without relying on traditional GNN explainer modules that often need training. Our approach produces concise explanation subgraphs with high fidelity and semantic clarity, outperforming prior methods. Logic demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
7 Ethical considerations
In this work, we have built methods for generating explanations for graph neural network predictions. We do not foresee any ethical issues from our study.
8 Limitations
We present some limitations of our work:
  • Our method projects GNN embeddings into the LLM embedding space and interleaves them with text to form hybrid prompts. However, batching multiple nodes leads to degraded performance and hallucinations, likely due to interference between soft prompts.
  • While Logic achieves high empirical fidelity, it does not provide formal guarantees of faithfulness to the GNN’s internal decision logic, as the explanation relies on the LLM’s interpretation of projected embeddings rather than directly tracing GNN computation paths.
  • Although our method incorporated a projector to align the embeddings of the GNN with the LLM space of token embeddings, the objectives we introduce for training the projector are not directly tailored for the task of explanation. More optimal objectives that adhere to the explanation task could bring substantial benefits.
Appendix A Additional Related Work
Here we provide additional related works that are on the integrated methods of LLMs and GNNs. The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017b); Hamilton et al. (2017). The approaches for integrating LLMs and GNNs primarily vary in the sequence of component application Ren et al. (2024); Chen et al. (2024b).
LLMs as prefix and postfix. First, there are methods that utilize LLMs at the prefix modules of their frameworks for task and feature representation Liu et al. (2024a); Fang et al. (2024); He et al. (2024a); Yang et al. (2025). This architecture provides greater flexibility, as various tasks can be specified in natural language with text-based graph representations, and LLM modules can remain fixed. On the other hand, the second class of methods uses GNNs as the prefix Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024), which is better suited for capturing a graph’s topology and its intricate relationships and uses LLMs for prediction. This approach also offers flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024).
Independent LLMs and joint LLM-GNNs. Additionally, there is a body of work that employs LLMs independently to tackle graph-related tasks Cheng et al. (2024); Fatemi et al. (2024); Li et al. (2024). Approaches following this paradigm benefit from the generalization capabilities of LLMs at both input and output levels, and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts. Finally, there is a group of works which employ Gnns and LLMs jointly through soft prompting, which necessitates alignments between their embeddings while enhancing LLMs reasoning by the Gnn captured graph structural features Liu et al. (2024b); Wang et al. (2024); Tian et al. (2023).
Explainers for GNNs. Several algorithms have been proposed in the literature to enhance the interpretability of GNNs. The majority of existing explainers concentrate on instance-level explanations. Instance-level (or local) explainers operate on individual graphs, identifying influential components—such as subgraphs—that most strongly impact the model’s prediction Ying et al. (2019b); Luo et al. (2020); Shan et al. (2021); Huang et al. (2022); Yuan et al. (2022); Lucic et al. (2022b); Tan et al. (2022b); Lin et al. (2021); Bajaj et al. (2021); Abrate and Bonchi (2021); Chhablani et al. (2024); Wellawatte et al. (2022); Verma et al. (2024); Lu et al. (2024). However, this localized perspective constrains their ability to uncover global patterns leveraged by GNNs across multiple graphs, as well as how these patterns integrate into a unified decision-making process. Research on global GNN explainers remains relatively limited Yuan et al. (2020); Huang et al. (2023); Xuanyuan et al. (2023); Azzolin et al. (2023); Armgaan et al. (2024); Fournier and Medya (2025). For instance, XGNN Yuan et al. (2020) and GNNInterpreter Wang and Shen (2023) adopt generative modeling approaches, producing graphs that most strongly correspond to a given class label. For more details, please refer to this survey on GNN explainers Kakkad et al. (2023b).
Table 4: Statistics of the graph datasets.
Dataset # Nodes # Edges # Categories # Features
Cora 2,708 5,429 7 2000
WikiCS 11701 148555 10 300
Amazon 1000 1397 47 2000
Liar 13293 32443 7 387
Appendix B Datasets
B.1 Datasets Details
We describe in detail the datasets that we used in our evaluation below. The basic statistics about these datasets are provided in Table 4.
  1. 1.
    Cora: It is a citation network, in which nodes represent computer science research papers, and each edge between two nodes represents a research paper citing another. The nodes are classified into one of seven categories. Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors. Each node’s feature is a 2000-dimensional bag-of-words representation of the keywords in the paper that it corresponds to.
  2. 2.
    WikiCS: It is a text-attributed graph dataset, derived from the Wikipedia platform, widely used for node classification tasks. The nodes correspond to Wikipedia page descriptions of different computer science topics, and edges between nodes represent hyperlinks from one article to another. Each node in the dataset belongs to one of 10 categories. The node classification task is to correctly predict a node’s label, and the edges are considered to be undirected. Each node’s feature vector is a GloVe embedding of size 300, computed based on the text associated with the nodes.
  3. 3.
    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about.To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set. Each node is embedded with a 384-dimensional BERT representation and augmented with a 3-dimensional one-hot vector indicating its type. Edges representing the original heterogeneous relations (spoken_by and about) are converted to undirected edges and unified into a single edge type in the homogeneous graph. Statement nodes retain their ground-truth labels for fake news classification (ranging from pants-fire to true), while speaker and topic nodes are assigned a dummy label. We also preserve a node-level string attribute—containing the raw statement, speaker name, or topic—for use in explanation and visualization.
  4. 4.
    Amazon-Product: It is a network with nodes representing different kinds of products on Amazon and edges connecting co-purchased products. The nodes belong to one of 47 different categories of products. We use only a subset of the Amazon-Product   dataset, consisting of the first 1000 products and their co-purchase edges, as the entire dataset is very large, and call it Amazon   in our work. The node classification task is to predict a product’s category based on its own as well as its neighbours’ features.
B.2 Relevance of datasets
Table 5: Dataset Descriptions and Relevance to Our Work
Dataset Description & Relevance Why It Fits Our Work
CORA Citation network of CS papers with bag-of-words features. Standard benchmark for node classification on text-rich graphs.
WIKICS Wikipedia CS articles linked by hyperlinks, with GloVe embeddings. Large, semantic-rich graph ideal for explanation evaluation.
LIAR Fake news detection graph combining statements, speakers, topics. Challenging, heterogeneous graph testing method adaptability.
AMAZON Product co-purchase network with 47 categories. Real-world, diverse e-commerce graph to test scalability.
The datasets used in our work are selected based on their relevance to graph explainability tasks, allowing us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features. The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake-news detection, and e-commerce, which reflect both real-world complexity and common challenges faced in explainability. The characteristics of each dataset make them suitable for evaluating the performance of our explanation method. Table 5 summarizes the key features of the datasets and explains their relevance.
Appendix C Additional Implementation Details
We discuss additional details related to the training of the base GNN.
C.1 GNN training
For all datasets, we implement a GCN with 3 convolutional layers, with the input and output channel dimensions chosen according to each dataset’s feature dimension and number of classes, respectively. For the hidden dimension, we choose either $64$ or $512$, depending on the nodes’ feature dimension. The model is trained with the Adam optimizer Kingma and Ba (2014) and a learning rate of 0.001 for 400 epochs.
C.2 GNN accuracy
We give the GNNs accuracy for the different folds in Table 6.
Table 6: Accuracy of GCN on the node classification task for the Cora, Wiki CS, Liar and Amazon products datasets.
Dataset Training Validation Testing
Cora 0.88 0.81 0.84
WikiCS 0.89 0.82 0.81
Amazon 0.87 0.73 0.76
Liar 0.42 0.30 0.29
Appendix D Prompt and generated responses examples
In this section we present a detailed prompt construction in D.1 and a complete example of such built prompt with its input in D.2.
D.1 Example of Full Prompt Construction for Logic
We have provided an example of the full prompt used in Logic below. The boldface texts denote where the embeddings are injected into the prompt. An example response is also illustrated in Figure 3 in the paper.
Prompt:
You are an expert research assistant. Your goal is to assess whether the classification of a target product is supported by its co-purchase neighborhood by analyzing Amazon product reviews and its co-purchase neighborhood to understand why it has been classified under a specific category. The categories to distinguish from are Home & Kitchen, Health & Personal Care, Beauty, etc.
Target Product ID: 0
Predicted Category: Clothing, Shoes & Jewelry
Target Product Embedding Representation:
[…]
Neighboring Products in the Co-purchase Network:
Each product below is described by keywords.
  • Product 129:
[…]
  • Product 130:
[…]
Instructions:
You are given a target product and a list of neighboring products, each described by keywords.
For each neighboring product:
  1. 1.
    Write one sentence summarizing the main topics or ideas captured in its keywords.
  2. 2.
    Clearly state whether this product supports the classification of the target product into category ‘Clothing, Shoes & Jewelry’.
Use the following format for each neighbor:
Product {ID}:
Summary: One sentence summary of the product’s keywords.
Support: YES or NO — Does this product support the classification into ‘Clothing, Shoes & Jewelry’?
Base your reasoning only on the keywords and proximity to the target product.
Start your analysis below:
D.2 Sample of the prompt and output for the Amazon-Product dataset
We provide a sample of the prompt and a sum up of the output we obtain (without the supporting sentence for each product) in Figure 3.
Figure 3: An illustration of the prompting template and LLM explanation of Logic for the Amazon dataset. The left block shows the input prompt to the LLM for a sample target node (target node id 3). The right block denotes the LLM response explaining the prediction of the GNN for the target node.
Appendix E Additional Experiments
We extend Table 1 of the paper with additional experiments for Amazon with TAGExplainer Pan et al. (2024), PGExplainer Luo et al. (2020) and with the GIN Xu et al. and GAT Veličković et al. (2017) architectures. The results are reported in Table 7.
Table 7: Performance comparison on the Products dataset using different GNN architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
GCN GAT GIN
Fidelity Size Fidelity Size Fidelity Size
Node 73.8% 1.00 73.2% 1.00 39.6% 1.00
Random 84.4% 7.24 84.2% 7.28 90.2% 7.25
GNNExplainer 95.5% 5.63 94.5% 5.615 88.0% 5.65
PGExplainer 87.2% 5.46 88.3% 5.76 87.2% 5.46
TAGExplainer 73.8% 1.37 71.8% 1.43 42.3% 1.27
LLM 86.5% 5.07 91.0% 5.00 86.5% 5.18
Our method 87.0% 1.3 81.5% 1.45 60.0% 1.52
Our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance for the GCN with 87.0% fidelity at just 1.3 in size. However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (60.0%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation. In contrast, GNNExplainer maintains the highest fidelity across all architectures but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes. The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise. These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structural-based models.
Table 8: Results on Products dataset using different pretrained LLMs (for pretrained GCN). Higher fidelity and lower size are better.
Llama 3.1 8B Instruct Mistral 7B Instruct v0.2 GPT-Neo 2.7B Phi-3 Mini 4k Instruct Pythia 2.8B
Fidelity Size Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 86.5% 5.07 88.0% 5.14 84.0% 2.33 88.0% 5.17 84.0% 2.34
Our method 87.0% 1.30 83.0% 1.63 80.0% 1.00 80.0% 1.00 80.0% 1.00
The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger. In contrast, our method consistently generates more compact explanations—typically around size 1, with a moderate decrease in fidelity of roughly 4–7 percentage points depending on the model. Notably, Llama 3.1 and Mistral backbones show strong fidelity for both methods, whereas smaller models like GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall. These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
Appendix F Definition of the qualitative metrics
We adapt the different dimensions of the framework in Cedro and Martens (2025) to the context of Cora to clarify their definitions, and provide the definition as given to the expert.
  1. 1.
    Understandability. Explanations should clearly convey why a given paper was assigned to a particular research topic.
  2. 2.
    Trustworthiness. Explanations should help users assess whether the model’s classification of a paper can be trusted.
  3. 3.
    Insightfulness. Explanations should reveal insights about the applications or connections that might play a role in the classification.
  4. 4.
    Satisfaction. Explanations should feel complete and meaningful to users evaluating the model’s behavior.
  5. 5.
    Confidence. Explanations should help users gain confidence in the correctness of the classification.
  6. 6.
    Convincingness. Explanations should be persuasive in justifying the model’s decision for a given paper.
  7. 7.
    Communicability. Explanations should be expressed in a way that aligns with the user’s background knowledge and expectations.
  8. 8.
    Usability. Explanations should support practical tasks such as interpreting predictions, or improving model performance.
Appendix G Experimental Setup for the human evaluation of M1 and M2 Scores
We clarify the experimental setup used to obtain the M1 and M2 scores reported in Table 2.
A) Data Preparation
We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained. For each article, we generated explanations using both LOGIC and GNNExplainer. Each expert was provided with:
  • The original article text and the GNN’s predicted label.
  • Two explanation subgraphs (one per method), anonymized and presented in random order known only to the experimenters.
  • For each subgraph, we listed the texts of the nodes it contains, sorted by increasing node index.
B) Evaluation Procedure
  • Before the evaluation, we provided experts with definitions of the evaluation metrics (as described in Appendix F) and asked them to consider how to apply each metric to content from the Cora dataset.
  • During the evaluation, each expert received the materials described above and independently rated each explanation on a scale from 1 to 5 for each metric.
C) Compilation of Results
The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Logic. These results are reported in Table 5.3.

### v3

@@
 Abstract
-Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs, which are common in domains such as citation networks, social platforms, and knowledge graphs.
+Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs (TAGs), which are common in domains such as citation networks, social platforms, and knowledge graphs.
 GNNs are not inherently interpretable and thus, many explanation methods have been proposed.
 However, existing explanation methods often struggle to generate interpretable, fine-grained rationales, especially when node attributes include rich natural language.
-In this work, we introduce Logic, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions.
-Logic projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure.
-This enables the LLM to reason about GNN internal representations and produce natural language explanations along with concise explanation subgraphs.
-Our experiments across four real-world TAG datasets demonstrate that Logic achieves a favorable trade-off between fidelity and sparsity, while significantly improving human-centric metrics such as insightfulness.
-Logic sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
+In this work, we introduce Gspell, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions.
+Gspell projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure.
+This enables the LLM to reason about GNN internal representations and to produce natural-language explanations, along with concise explanation subgraphs.
+Our experiments across real-world TAG datasets demonstrate that Gspell achieves a favorable trade-off between fidelity and sparsity, while improving human-centric metrics such as insightfulness.
+Gspell sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
 From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context
@@
 (2021); Yuan et al.
+(2021); Armgaan et al.
 (2024), or offer counterfactual scenarios Lucic et al.
 (2022a); Tan et al.
-(2022a); Chen et al.
-(2022).
+(2022a); Verma et al.
+(2024); Huang et al.
+(2023); Fournier and Medya (2025).
 However, generating explanations for GNNs is inherently challenging due to the combinatorial nature of graph data and the joint influence of node attributes and edge connections.
 These complexities make GNN explainability a non-trivial problem and have led to the development of a wide array of explanation techniques Kakkad et al.
-(2023a); Guo et al.
+(2023); Guo et al.
 (2023).
-1.1 Need for LLM-based GNN explainers
-There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text attribute graphs (TAGs).
+Need for LLM-based GNN explainers.
+There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text-attributed graphs (TAGs).
 Essentially, these methods leverage LLMs in different architectural orders to 1) attain rich representations of the input TAGs, task descriptions, and task-specific few-shot examples for the GNN predictors Liu et al.
@@
 (2024).
-However, there is a lack of focus on using LLMs to explain GNNs predictions and clarify the reasoning behind these black-box models.
+However, there is a lack of focus on using LLMs to explain GNN predictions and clarify the reasoning behind these black-box models.
 Traditional GNN explainers aim to identify influential subgraphs or feature subsets that contribute to predictions Ying et al.
@@
 (2020).
-While effective for graphs with simple node features, these methods struggle in TAGs, where node information is embedded in natural language.
-Moreover, often these explanations are in the form of subgraphs that are not human-tangible for domain users.
+While effective for graphs with simple node features, these methods perform poorly on TAGs, where node information is expressed in natural language.
+Moreover, these explanations are often presented as subgraphs that are not human-interpretable to users.
 LLMs, on the other hand, excel at reasoning over textual content and can generate coherent, human-interpretable rationales.
-Thus, their integration into GNNs explanation pipelines can improve the interpretability of the GNN models outputs on TAGs.
-1.2 Limitations of Existing LLM-based GNN Explainers
-The existing frameworks (please see Sec.
-2) that use LLMs to build GNN explanations have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainers’ output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or extensive training/fine-tuning of the GNN explainer.
+Thus, their integration into GNN explanation pipelines can improve the interpretability of the GNN outputs on TAGs.
+Limitations of Existing LLM-based GNN Explainers.
+The existing frameworks (detailed in Sec.
+2) employing LLMs for explaining GNN have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainer’s output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or training/fine-tuning of the GNN explainer.
 (2) Missed GNN internals: Existing approaches fail to directly leverage the rich internal representations of the GNN, making the explanations generic or unfaithful to the internal working of the GNNs.
-(3) Bias from suboptimal explainers: Involving a pre-existing GNN explainer can bias the reasoning and judgment of the LLM.
-This is critical when the GNN explainers are suboptimal and they signal noisy information to LLM, while the LLM can potentially better infer and reason the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
-1.3 Our Contributions
-In this work, we present Logic (Language mOdel-guided GNN Interpretation with Context), a lightweight, post-hoc explanation framework for text-attributed graphs (TAGs) that seamlessly integrates the representational power of GNNs with the reasoning capabilities of LLMs.
+(3) Bias from suboptimal explainers: Invoking an external GNN explainer can bias the reasoning and judgment of the LLM.
+This is critical when the GNN explainers are suboptimal, and they signal noisy information to LLM, while the LLM can potentially better infer the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
+Our Contributions.
+In this work, we present Gspell (GNN Soft Prompted Explanation with LLMs), a lightweight, post-hoc explanation framework for TAGs that integrates the representational power of GNNs with the reasoning capability of LLMs.
 Our contributions are as follows:
-  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by directly using LLMs as interpreters of GNN behavior.
-Logic generates fine-grained, natural language explanations without relying on handcrafted templates or perturbation-based saliency masks, reducing bias from external explainers.
-  • Soft Prompt Integration: Logic aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector.
+  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by using LLMs as direct interpreters of model behavior.
+Gspell generates fine-grained, natural language explanations (NLEs) unreliant on handcrafted templates or perturbation-based saliency masks, reducing external explainers’ bias.
+  • Soft Prompt Integration: Gspell aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector.
 This projection enables GNN features to be treated as soft prompts in a hybrid prompt that blends structural and textual cues, allowing the LLM to reason over the graph’s latent space.
-  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Logic produces interpretable subgraphs based on node-level binary support decisions derived from natural language explanations.
-  • Training-Agnostic and Plug-and-Play Deployment: Logic does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining.
+  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Gspell produces interpretable NLE subgraphs of node-level support decisions.
+  • Training-Agnostic and Plug-and-Play Deployment: Gspell does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining.
 This makes it adaptable for real-world applications.
@@
 We describe the related studies on the integrated LLM and GNN explainer methods.
-The explanation capability of LLMs has been applied to graph problems in task-oriented contexts such as explainable recommendation systems Li et al.
+The explanation capability of LLMs has been applied to graph problems in task-oriented contexts Li et al.
 (2025); Ma et al.
 (2024).
-There has been attempts to guide GNNs using the explanations generated by LLMs, for example, in He et al.
-(2024a), an LLM initially generates node predictions and textual explanations on the reasoning behind the predictions.
-These explanations then guide a second LLM to produce new node features, which augment the original text features as input for a GNN predictor.
+There have been attempts to guide GNNs using explanations generated by LLMs; for example, in He et al.
+(2024a), an LLM generates node predictions and textual explanations of the reasoning behind them.
 LLM explanation to enhance explainers.
-LLMs have been utilized to enhance the explanations of pre-existing GNN explainers, for example Zhang et al.
-(2024) uses LLMs to evaluate explanations of an explainer, its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias.
+LLMs have been utilized to enhance the explanations of another GNN explainers, for example Zhang et al.
+(2024) uses LLMs to evaluate explanations of an explainer and its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias.
 Similarly, to improve the understandability of explanations, Pan et al.
-(2024) first trains a pseudo-label generator LLM that takes explanations from a pre-existing GNN explainers and generates pseudo-labels for them.
+(2024) first trains a pseudo-label generator LLM that takes explanations from an external GNN explainers and generates pseudo-labels for them.
 The generator is tuned in an expert iteration procedure with tailored objectives for faithfulness and brevity, and the optimized generator generates pseudo-labels to fine-tune an explainer LLM to build end-to-end models.
 LLM explanation to enhance understanding.
-One advantage of LLMs is that they can generate natural language explanations of GNNs, which are more interpretable for humans.
+One advantage of LLMs is that they can generate NLE for GNNs, which are more human-interpretable.
 He et al.
-(2024b) also proposes a method for generating counterfactual graph explanations for GNNs, which uses a counterfactual autoencoder (CA)—composed of a contrastively pretrained text encoder based on LLM-generated text pairs and a graph decoder—to construct counterfactual graph topologies by LLM-generated counterfactual text pairs.
-To address hallucination, the framework incorporates a dynamic feedback mechanism to refine the generated counterfactual text pairs by prompting the LLM back again first-round generated counterfactuals.
-Concurrent to our work Cedro and Martens (2025) suggests a narrative (story-like) explanation for a GNN’s predictions.
-Given an LLM, the dataset description, and a GNN model’s predictions, along with the explanation generated by a GNN explainer in the form of subgraphs and feature importance matrices, they ask the LLM to narrate the relationship between the neighboring nodes, the edges connecting them, and the node’s features but fall short of having the LLM as a method to generate the explanations.
-Our work aims at addressing these limitations and producing a GNN internal representation-aware LLM explanation.
+(2024b) proposes a method for generating counterfactual explanations, which uses an autoencoder to construct counterfactual graph topologies from LLM-generated counterfactual text pairs.
+To mitigate hallucinations, it employs a dynamic feedback mechanism that prompts the LLM to refine its initial outputs.
+Lately, Cedro and Martens (2025) proposed prompting an LLM to narrate the explanation subgraphs and the feature importance matrices generated by a GNN explainer, which describe the relationships between neighboring nodes.
+Our work produces representation-aware LLM explanations without an external explainer.
 We present additional related work in Appendix A.
@@
 Definition 1 (Text-Attributed Graph (TAG)).
-A Text-Attributed Graph (TAG) is defined as a graph
+A TAG is defined as a graph
 $G=(V,E,T)$
@@
 Each node $v\in V$ is thus associated with a text document $T(v)$, and the graph structure $E$ represents semantic, relational, or structural links between nodes.
-TAGs enable learning over both the textual content and the topology of the graph.
+TAGs enable learning from both the textual content and the graph’s topology.
 In this work, we focus on the node classification task.
 Node Classification.
-Given a dataset $\mathcal{G}$ and a graph $G=(V,E,X)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$.
-Our method aims at explaining the node classification predictions of a GNN with local subgraphs, i.e, small subgraphs around the node staged for classification Ying et al. (2019a). More formally:
+Given a dataset $\mathcal{G}$ and a graph $G=(V,E,T)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$, and $x_{v}=f_{e}(T(v))$, where $f_{e}$ is a text encoder.
+Our method aims to explain the node classification predictions of a GNN using local subgraphs, i.e., small subgraphs around the node being classified Ying et al. (2019a).
 Definition 2 (Local Factual Explainer).
 Given a GNN $\Phi$ and a graph $G=(V,E,T)\in\mathcal{G}$, a local factual explainer $\Psi$ is a mapping from $(\mathcal{V},\mathcal{G})$ to $\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ associated to $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ verifying:
-$\Phi(v,\Psi(v,G))=\Phi(v,G)\;\;\forall G\in\mathcal{G},\forall v\in V\cap\mathcal{V}^{\prime}$ (1)
+$\Phi(x_{v},\Psi(v,G))=\Phi(x_{v},G)\;\forall G\in\mathcal{G},v\in V\cap\mathcal{V}^{\prime}$
 where $\mathcal{S}_{v}(G)\in\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ denotes the set of all subgraphs of $G$ that contain $v$.
 $\mathcal{V}^{\prime}$ corresponds to the set of nodes over all graphs for which the equation is verified, and in this case, the explanation $\Psi(v,G)$ is called faithful.
-Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ verifies Equation 1 for every node.
-However, such explainers are rarely useful in practice, as they do not isolate concise or interpretable subgraphs.
-We therefore evaluate explanations based on their size and fidelity, as formally defined in Section 5.1.
+Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ (identity function) verifies Equation 1 for every node.
+However, such explainers are rarely useful in practice because they do not produce concise, interpretable subgraphs.
+Therefore, the explanation subgraph size is an important metric, as formally defined in Section 5.1.
 Although Definition 2 necessitates faithfulness of the explanation graphs and the existing methods that build explanations for GNNs mostly follow this definition, there are two major problems:
   • It does not necessarily generate concise explanations.
-Generally, a good factual explanation subgraph is a sparse one Kakkad et al.
-(2023a); Yuan et al.
-(2023) which has only the most important features of the graph, i.e., nodes, edges, and node features.
-This property makes the graph more comprehensible for users.
-  • More importantly, this problem definition does not account for human-level understanding of the explanations.
-For example, retrieving a subgraph as a common form of explanations Ying et al.
+Generally, a good factual explanation subgraph is sparse Kakkad et al.
+(2023); Yuan et al.
+(2023), containing only the most important graph elements: nodes, edges, and node features.
+  • More importantly, this problem definition does not account for human-level interpretability.
+For example, retrieving an explanation subgraph is common Ying et al.
 (2019a); Luo et al.
-(2020) needs additional information for the domain-experts to understand the results and make them useful in practice.
-Nevertheless, these two desirable criteria are indeed compatible and aligned with each other as well as the objective of explainable AI (XAI), which attempts to make black-box AI models more interpretable for humans.
-Therefore, our goal is to design a method to enhance the understandability of the explanations while making them coherent and concise.
-Figure 1: Illustration of Logic’s framework.
+(2020), but it lacks sufficient information for understanding, which limits its usefulness in practice.
+These criteria align with the goals of Explainable AI (XAI). Consequently, we aim to develop a method that enhances interpretability while maintaining conciseness.
+Figure 1: Illustration of Gspell’s framework.
 First, the projector is trained to align GNN node embeddings with the LLM’s embedding space.
 Next, hybrid prompts are constructed by interleaving projected embeddings (as soft prompts) with natural language tokens.
-These prompts are then fed into the LLM to produce natural language explanations, which are converted into explanation subgraphs.
-4 Our Proposed Method: Logic
-To explain the predictions of a trained GNN using an LLM, our method projects node embeddings from the GNN latent space into the LLM embedding space, allowing the LLM to reason about the GNN’s predictions.
-Rather than relying solely on textual prompts, our approach injects GNN embeddings as soft prompts interleaved with natural language tokens.
-This enables the LLM to generate context-aware explanations grounded in both the model’s internal representations and human-interpretable features.
-We present our method in three parts: first, the training of the projector; second, the construction of the hybrid prompt; and third, the generation of the explanation subgraph.
-An overview of our approach is shown in Figure 1.
+These prompts are then fed to the LLM to produce natural language explanations, which are converted into explanation subgraphs.
+4 Our Proposed Method: Gspell
+To explain a trained GNN’s predictions, our method projects node embeddings from the GNN’s latent space into the LLM’s embedding space.
+By injecting these embeddings as soft prompts interleaved with natural language, we enable the LLM to generate context-aware explanations grounded in the model’s internal representations.
+Our approach consists of three parts: training the projector, constructing the hybrid prompt, and generating the explanation subgraph (Figure 1).
 4.1 Projection from the GNN
-In order for the LLM to explain the GNN model, we aim to build a framework where the LLM is able to understand how the GNN sees each node of the computation tree when classifying a particular node.
-To that effect, we introduce a projector that maps nodes’ GNN embedding into tokens.
-Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer.
-The projector aims at both aligning its input with the GNN embeddings and grounding its output in the natural language context of the nodes, hence considering two different levels of embeddings.
+To help the LLM interpret the GNN, we developed a framework that aligns the LLM with the GNN’s view of the computation tree.
+Specifically, we use a projector to map GNN node embeddings directly into the LLM’s token space.
+Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E,T)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer.
+The projector aligns GNN embeddings with natural language context, effectively bridging two distinct embedding spaces.
 Generally speaking, this approach has similarities with other multi-modal alignment frameworks.
@@
 Similarly, our designed projector learns to map GNN embeddings into soft prompt tokens that serve as input to an LLM, effectively bridging the GNN’s latent space and the LLM’s token-level semantics.
-Formally, we define a projector as a function:
-$\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$
-where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings.
+Formally, we define a projector as a function $\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$, where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings.
 We write $\Pi(f_{\Phi}(v))=Z_{v}\in\mathbb{R}^{k\times h}$.
-The projector $\Pi$ is trained to optimize two losses:
+$\Pi$ optimizes two losses:
   1. 1.
     Context alignment loss: This encourages the average soft token representation to align with the LLM embedding of the natural language text associated with the node:
-    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\left\langle\bar{Z}_{v},\;LLM(v)\right\rangle\right]$
-    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output (normalized), and $LLM(v)\in\mathbb{R}^{k}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space.
-The inner product corresponds to cosine similarity.
+    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\text{cos}\left(\bar{Z}_{v},\;LLM(v)\right)\right]$
+    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output, $\text{cos}(.,.)$ denotes cosine similarity, and $LLM(v)\in\mathbb{R}^{h}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space.
   2. 2.
@@
     where
-    $p_{vu}^{\Phi}=\frac{\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(u)\right\rangle/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(w)\right\rangle/\tau\right)}$
-    $p_{vu}^{\Pi}=\frac{\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{u}\right\rangle\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{w}\right\rangle\right)}$
-    and $\tau>0$ is a temperature hyperparameter. Both similarity scores are computed using cosine similarity between normalized vectors.
-    The overall training objective becomes:
-    $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$
-    where $1\geq\beta\geq 0$ controls the balance between the two losses.
+    $p_{vu}^{\Phi}=\frac{\exp\left(\text{cos}\left(f_{\Phi}(v),f_{\Phi}(u)\right)/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\text{cos}\left(f_{\Phi}(v),f_{\Phi}(w)\right)/\tau\right)}$
+    $p_{vu}^{\Pi}=\frac{\exp\left(\text{cos}\left(\bar{Z}_{v},\bar{Z}_{u}\right)\right)}{\sum_{w\in\mathcal{V}}\exp\left(\text{cos}\left(\bar{Z}_{v},\bar{Z}_{w}\right)\right)}$
+    and $\tau>0$ is a temperature hyperparameter.
+    The overall training objective becomes: $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$ where $1\geq\beta\geq 0$.
 4.2 Prompt Construction
 Given a node $u$, the projector outputs a matrix $\Pi(f_{\Phi}(u))=Z_{u}\in\mathbb{R}^{k\times h}$, which we refer to as a soft prompt embedding.
-This matrix is injected directly into the LLM’s embedding space and interleaved with natural language token embeddings.
+This matrix is directly injected into the LLM’s embedding space and interleaved with natural language token embeddings.
 For a node $v$ that we wish to explain, we consider its computation tree for GNN $\Phi$, denoted by $\mathcal{T}^{\phi}_{v}$.
-$\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$, the number of layers of $\Phi$, composed of all paths of length $L$ starting at $v$, with repetitions, concatenated to the root $v$.
-We construct the LLM input as a sequence of embeddings consisting of:
-  • Standard token embeddings for the system prompt and user instruction.
-  • A segment labeled “Target Paper Embedding Representation,” where we insert the soft prompt matrix $Z_{v}$ for the target node, bracketed by textual markers.
-  • An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is encoded by interleaved natural language markers and its own soft-prompt embeddings.
-  • Final instructions text prompting the model to provide a summary and binary support decision for each computation tree node based on the embeddings.
-The LLM sees the embedded GNN features as native tokens, enabling reasoning over both GNN’s representation and natural language inputs. We give an example of the resulting prompt in D.
+$\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$ (the number of layers of $\Phi$) composed of all paths of length $L$ starting at $v$, concatenated to the root $v$.
+We construct the LLM input as a sequence of embeddings consisting of: (i) Standard token embeddings for the system prompt and user instruction; (ii) The target node’s $k$-token embedding $Z_{v}$—generated after enclosing the node’s text descriptor by text markers; (iii) An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is enclosed by text markers and its own soft-prompt embeddings; and (iv) Final query instructions.
+By treating the GNN embeddings as native tokens, the LLM reasons across both GNN representations and textual features (see Figure 2 and Appendix D.2).
 4.3 Explanation Generation
-To produce the explanations, the LLM first predicts whether each node in the computation tree supports or does not support the target node’s classification.
-This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, (iii) Neutral nodes, not assigned to either category.
-The explanation subgraph is then formed by including all supporting nodes and a controlled number of neutral nodes.
-The inclusion of hallucinated nodes is proportional to the number of supporting nodes relative to the total size of the computation tree, as described in our size constraint formulation.
-For each node $v$, the LLM generates a natural language explanation $E_{v}$ considering the influence of nodes of $\mathcal{T}^{\Phi}_{v}$ with regards to the classification of node $v$.
-We define a label attribution function based on the parsing of $E_{v}$:
-$\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$
-such that for each node $u\in\mathcal{T}^{\phi}_{v}$:
+First, the LLM predicts whether each node in the computation tree supports or does not support the target node’s classification.
+This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, and (iii) Neutral nodes, which do not appear in either category.
+Formally, the LLM generates a natural language explanation $E_{v}$ based on $\mathcal{T}^{\Phi}_{v}$ and the classification of $v$.
+Given the parsing of $E_{v}$ we apply a label attribution function $\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$ for each node $u\in\mathcal{T}^{\phi}_{v}$ such that:
   • $\chi(u)=1$ if $u$ is cited in $E_{v}$ as supporting the classification of $v$,
@@
 $S^{0}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=0\}$
-To address uncertainty and enforce conciseness, we introduce a refinement function $\rho:2^{\mathcal{T}^{\phi}_{v}}\to 2^{\mathcal{T}^{\phi}_{v}}$ such that the final explanation subgraph is:
-$S_{v}=\rho(S^{+}_{v},S^{-}_{v},S^{0}_{v})$
-[…]
+The final explanation subgraph is $S_{v}=S^{+}_{v}$.
 Mitigating Hallucination.
@@
 We address this issue through two complementary mechanisms within our framework:
-  • Prompt templating with constrained node sets: To reduce hallucinations, we ensure that our prompts contain a fixed set of nodes.
-We explicitly instruct the LLM to generate explanations only referencing these nodes, limiting the possibility of hallucinating unsupported nodes.
-  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes are filtered out in a post-processing step.
-This filtering process ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
+  • Prompt templating with constrained node sets: To reduce hallucinations, we partition $\mathcal{T}^{\phi}_{v}$ into fixed-size batches and sequentially prompt them to the LLM, instructing it to generate explanations that reference only the nodes in each batch.
+  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes—the nodes absent in $\mathcal{T}^{\phi}_{v}$—are filtered out in a post-processing step.
+This filtering ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
 These strategies work together to minimize hallucination and enhance the accuracy and reliability of the generated explanations.
+4.4 Faithfulness to Gnn Predictions
+Beyond constraining hallucinations, we also aim to ensure that the generated explanations remain faithful to the Gnn’s predictions.
+Accounting for context aligns the GNN embeddings with those of the LLM; however, this alignment alone does not guarantee that the GNN’s decisions are faithful to the LLM’s.
+We address faithfulness in two ways.
+First, we explicitly include the Gnn’s predicted label for the target node in the input prompt (Sec.
+4.2), so that the LLM conditions its reasoning on the model’s actual decision.
+This provides the LLM with a clear natural-language interpretation of the prediction target.
+Second, we introduce Gspell+, which directly optimizes soft prompts to ensure faithfulness to the Gnn’s predictions.
+Assume the Gnn is a predictive model $\Phi=h\circ g$ where $g:G\rightarrow\mathcal{E}$ is a message-passing encoder and $h:\mathcal{E}\rightarrow Y$ is a decoder s.t.
+$\mathcal{E}\in\mathbb{R}^{m}$ and $Y\in\mathbb{R}^{C}$ are the latent and output spaces, and $C$ is the number of classes.
+If $\Phi$ is well optimized, the conditional entropy $H(Y\mid\mathcal{E})$ is low due to the pretraining stage, which is equivalent to increasing the mutual information $I(Y;\mathcal{E})$.
+Therefore, a good projector $\Pi$ has to not only align the Gnn embeddings with the LLM embeddings, but also have the projections share high mutual information with the output of the Gnn.
+Along with the previous alignment objectives $\mathcal{L}_{\text{context}}$ and $\mathcal{L}_{\text{contrast}}$, we further consider increasing the mutual information $I(\hat{y}_{v};Z_{v})$, where $\hat{y}_{v}=h(g(v))\in\mathbb{R}^{C}$ by minimizing the InfoNCE loss van den Oord et al.
+(2019).
+Since $Z_{v}$ and $\hat{y}_{v}$ do not live in the same space, we use another projector $\Pi^{\prime}:\mathbb{R}^{k\times h}\rightarrow\mathbb{R}^{C}$ to back-propagate through the original projector $\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$ the faithfulness to the prediction.
+Therefore, writing $Z^{\prime}_{v}=\Pi^{\prime}(Z_{v})$, the final objective for Gspell+ becomes:
+$\mathcal{L}=\beta_{1}\;\mathcal{L}_{\text{context}}+\beta_{2}\mathcal{L}_{\text{contrast}}+\beta_{3}\mathcal{L}_{\text{InfoNCE}}\text{ where}$
+$\mathcal{L}_{\text{InfoNCE}}=$
+$-\mathbb{E}\left[\log\frac{\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{v})}{\tau}\right)}{\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{v})}{\tau}\right)+\underset{u\ \neq v}{\sum}\exp\left(\frac{\text{cos}(Z^{\prime}_{v},\hat{y}_{u})}{\tau}\right)}\right]$
+Amazon Cora Liar WikiCS
+Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
+GNNExplainer 94.5 $\pm$ 1.5 5.71 $\pm$ 0.01 96.5 $\pm$ 0.6 17.91 $\pm$ 0.01 100.0 $\pm$ 0.00 474.01 $\pm$ 0.00 99.0 $\pm$ 0.0 1930.89 $\pm$ 0.0
+PGExplainer 73.8 $\pm$ 5.6 4.52 $\pm$ 0.13 82.0 $\pm$ 1.0 8.96 $\pm$ 0.40 100.0 $\pm$ 0.0 *1.0 90.6 $\pm$ 0.0 295.89 $\pm$ 0.0
+TAGExplainer 65.1 $\pm$ 4.4 1.03 $\pm$ 0.00 78.5 $\pm$ 0.9 1.06 $\pm$ 0.00 100.0 $\pm$ 0.0 1.31 $\pm$ 0.00 76.4 $\pm$ 1.1 1.51 $\pm$ 0.00
+[…]
+Random 91.0 $\pm$ 1.1 7.28 $\pm$ 0.02 93.7 $\pm$ 0.5 18.67 $\pm$ 0.01 100.0 $\pm$ 0.0 469.36 $\pm$ 0.03 93.3 $\pm$ 0.0 994.76 $\pm$ 0.0
+LLM 86.0 $\pm$ 1.2 4.90 $\pm$ 0.23 94.5 $\pm$ 1.0 3.27 $\pm$ 0.31 100.0 $\pm$ 0.0 2.22 $\pm$ 0.04 91.4 $\pm$ 1.0 21.19 $\pm$ 0.30
+Gspell+ 92.0 $\pm$ 1.8 4.24 $\pm$ 0.50 85.8 $\pm$ 1.8 2.10 $\pm$ 0.11 100.0 $\pm$ 0.0 1.10 $\pm$ 0.08 86.3 $\pm$ 1.7 12.44 $\pm$ 0.27
+Gspell 91.1 $\pm$ 2.3 4.25 $\pm$ 0.54 86.5 $\pm$ 0.0 2.07 $\pm$ 0.07 100.0 $\pm$ 0.0 1.11 $\pm$ 0.08 86.5 $\pm$ 0.0 12.80 $\pm$ 0.0
 Table 1: Performance comparison with baseline models on the node classification task.
@@
 *The explainer returns all-zero masks on Liar, we consider its output to be the central node only.
-Cora WikiCS Liar Amazon
-Fidelity Size Fidelity Size Fidelity Size Fidelity Size
-Node 73.6% 1 77.8% 1 95.0% 1 73.8% 1
-Random (50%) 93.5% 18.1 88.9% 19.5 95.0% 19.0 84.4% 7.24
-Random (25%) 79.3% 10.0 86.9% 10.0 95.0% 18.13 73.3% 4.03
-GNNExplainer 92.0% 18.38 95.4% 9.7 100.0% 12.4 95.5% 5.63
-PGExplainer 86.5% 16.2 86.2% 21.2 95.0%* 1.00* 89.0% 5.73
-TAGExplainer 60.0% 1.30 80.0% 1.20 95.0%* 1.00* 73.8% 1.52
-LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.07
-Our method 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 1.30
+LLM Prompt You are analyzing Amazon product reviews and their co-purchase neighborhoods to understand why it has been classified under a specific category.
+Your task is to determine which of the neighboring products does (not) support the classification of the target product.
+Target Product ID: 160 Predicted Category: Sports & Outdoors
+Product Embedding Representation:
+<<BEGIN TARGET KEYWORDS>> PROJECTED EMBEDDING <<END TARGET KEYWORDS>>
+Neighboring Products:
+Each Product below is described by keywords.
+Product 3: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 4: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 5: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 10: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 12: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+LLM Response
+P3 Summary: This product seems to be a fitness accessory, possibly a yoga mat or a resistance band, given its mention of “workout” and “stretching”. YES
+P4 Summary: This product appears to be a nutrition supplement, possibly a protein powder or a vitamin, given its mention of “muscle” and “energy”. NO
+P5 Summary: This product seems to be a fitness tracker or a smartwatch, given its mention of “fitness” and “activity tracking”. YES
+P10 Summary: This product seems to be a health and wellness product, possibly a meditation app or a mindfulness journal, given its mention of “relax” and “meditate”. NO
+P12 Summary: This product appears to be a sports accessory, possibly a water bottle or a towel, given its mention of “hydrate” and “dry off”. YES
+Instructions For each neighboring product: 1.
+Summarize the similarity between supporting products in one sentence.
+2.
+Clearly state whether this product supports the classification of the Target Product into category Sports & Outdoors.
+Format:
+Product <ID>:
+Summary: <One sentence>.
+Support: YES or NO — Does this product support classification into Sports & Outdoors?
+Base reasoning only on the keywords and proximity to the target product.
+Figure 2: Left: prompt with category and embeddings highlighted. Right: full model response with summaries in the middle and YES/NO verdicts aligned on the right.
 5 Experiments
-Our method, Logic, acts as a light-weight framework to generate explanations of Gnns especially for TAGs.
-We demonstrate the effectiveness of Logic in various settings.
-Our code is publicly available at: https://github.com/pbaghershahi/LOGIC.git.
-5.1 Evaluation metrics
-We use the following metrics in the experiments:
-1) Fidelity: Fidelity captures how well an explainable model reproduces or approximates the GNN model logic through its prediction.
-Mathematically, the fidelity of the explainer $\Psi$ of GNN $\Phi$ on dataset $\mathcal{G}$ over $\mathcal{V}$ nodes is:
-$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathds{1}(\Phi(v,\Psi(v,G))=\Phi(v,G))$
-2) Explanation size: The explanation size measures the complexity or compactness of the explanation subgraph produced by the explainer $\Psi$.
-Specifically, it captures the number of nodes in the explanation for a given prediction.
-Smaller explanation sizes are generally preferred, as they are easier to interpret.
-The explanation size is:
-$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|\Psi(v,G)|$
+We demonstrate the effectiveness of Gspell in various settings. Our code is publicly available at: https://github.com/pbaghershahi/GSPELL.git.
 5.1 Experimental setup
+Evaluation Metrics.
+Having an explainer $\Psi$ of GNN $\Phi$, a dataset $\mathcal{G}$ of $\mathcal{V}$ nodes, and denoting explanation subgraph by $S_{v}=\Psi(v,G)$, we use the following metrics in the experiments:
+1) Fidelity:
+Fidelity captures how well an explainable model reproduces the GNN model’s logic in its predictions. Mathematically, we have:
+$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathbb{I}[\Phi(v,S_{v})=\Phi(v,G)]$
+where $\mathbb{I}[\cdot]$ is the indicator function.
+2) Size:
+Smaller explanation sizes are easier to interpret. The explanation size measures the compactness of the explanation subgraph, i.e., the number of nodes in the explanation, as:
+$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|S_{v}|$
+For a more comprehensive evaluation, we adopt the perturbation-based framework of recent natural language explanation literature Atanasova et al. (2023); Siegel et al. (2024).
+3) Comprehensiveness:
+Comprehensiveness measures the drop in prediction accuracy after removing the explanation subgraph.
+Let $G\setminus S_{v}$ denote the graph obtained by removing the nodes in $S_{v}$.
+Then, we define comprehensiveness as:
+$Comp.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(x_{v},G)=y_{v}\right]$-
+$\mathbb{I}\!\left[\Phi\!\left(x_{v},G\setminus S_{v}\right)=y_{v}\right],$
+4) Alignment:
+Alignment evaluates whether the nodes identified as opposing the prediction, the set $S_{v}^{-}$, are indeed unsupportive.
+We define alignment as the drop in prediction accuracy after removing these nodes:
+$Align.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(v,G)=y_{v}\right]$-
+$\mathbb{I}\!\left[\Phi\!\left(v,G\setminus S_{v}^{-}\right)=y_{v}\right].$
+5) Random Baseline:
+The random baseline measures the drop in prediction accuracy after removing a random node set, the same size as the explanation subgraph.
+Let $S_{v}^{\mathrm{rand}}\subseteq T_{v}^{\phi}$ be a uniformly sampled subset such that $|S_{v}^{\mathrm{rand}}|=|S_{v}|$, we define:
+$Rand.=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},\,v\in V}\mathbb{I}\!\left[\Phi(v,G)=y_{v}\right]$-
+$\mathbb{I}\!\left[\Phi\!\left(v,G\setminus S_{v}^{\mathrm{rand}}\right)=y_{v}\right].$
+Overall, higher values are preferred for Fidelity and Comprehensiveness, but lower values are preferred for Size and Alignment.
+The Random Baseline serves as a reference, where we expect Comprehensiveness to be significantly higher than the Random Baseline.
 Datasets.
 We use real-world TAG datasets to evaluate the performance of our method: Cora Sen et al.
-(2008) - a co-citation network of research papers, WikiCS Mernyei and Cangea (2022) - a network of related Wikipedia articles, the Amazon-Product Feng et al.
-(2024) dataset - an E-commerce product co-purchase network, and the Liar Wang (2017) dataset - a fake news detection knowledge-graph.
-Additional details about the processing and relevance of the datasets are provided in Appendix B.
+(2008), WikiCS Mernyei and Cangea (2022), Amazon-Product Feng et al.
+(2024), and Liar Wang (2017).
+Additional details of the datasets are provided in Appendix B.
 Models.
-We consider a GCN Kipf and Welling (2017a) as the base GNN model for the node classification task for all datasets.
-We train a model consisting of three convolutional layers, with the last one being the classification layer.
-The training/validation/testing split is $60\%/20\%/20\%$ for all the datasets.
-We give additional information about the GCN training in Appendix C.
-We use Meta’s Llama-3.1-8B-Instruct from the HuggingFace library as the frozen LLM in our experiments.
-We run all experiments on NVIDIA-GeForce-RTX-3090 GPUs, 24GB RAM.
-Figure 2: A visualization of the explanations on the Amazon dataset.
-Left shows the explanations subgraph generated by GNNExplainer.
-Right shows the explanation generated by Logic.
-The node marked by a bold circle (product #7) is the node of interest for the explanation.
-Nodes with the same color are from the same category.
-Although both models have similar quantitative performance (in terms of size and fidelity), the LLM-generated explanations are textually clarified, making them more human-understandable.
+We consider a GCN Kipf and Welling (2017b) as the base GNN model for the node classification task, and the LLM model is Meta’s Llama-3.1-8B-Instruct, which remains frozen in our experiments.
+Full experimental details are provided in Appendix D.
+We provide experiments with other LLM models in Appendix C.2 and for different GNN architectures in Appendix C.1.
+Dimension Cora Amazon-Product
+$M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta$ $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta$
+Understandability $2.9_{\scriptstyle\pm 1.23}$ $3.2_{\scriptstyle\pm 1.45}$ $0.3$ $3.33_{\scriptstyle\pm 1.46}$ $3.25_{\scriptstyle\pm 1.33}$ 0.08-0.08
+Trustworthiness $2.7_{\scriptstyle\pm 1.10}$ $3.1_{\scriptstyle\pm 1.06}$ $0.4$ $3.30_{\scriptstyle\pm 1.56}$ $3.05_{\scriptstyle\pm 1.41}$ 0.25-0.25
+Insightfulness $3.3_{\scriptstyle\pm 0.74}$ $3.1_{\scriptstyle\pm 0.95}$ 0.2-0.2 $2.95_{\scriptstyle\pm 1.34}$ $3.25_{\scriptstyle\pm 1.17}$ $0.30$
+Satisfaction $2.2_{\scriptstyle\pm 1.16}$ $2.7_{\scriptstyle\pm 0.92}$ $0.5$ $3.08_{\scriptstyle\pm 1.46}$ $3.05_{\scriptstyle\pm 1.36}$ 0.03-0.03
+Confidence $0_{\scriptstyle\pm 1.49} 3.3$$3_{\scriptstyle\pm 1.49}$ $0.3$ $3.13_{\scriptstyle\pm 1.47}$ $3.03_{\scriptstyle\pm 1.33}$ 0.10-0.10
+Convincingness $3.3_{\scriptstyle\pm 1.57}$ $3.7_{\scriptstyle\pm 1.64}$ $0.4$ $3.08_{\scriptstyle\pm 1.46}$ $2.93_{\scriptstyle\pm 1.49}$ 0.15-0.15
+Communicability $3.0_{\scriptstyle\pm 1.60}$ $3.1_{\scriptstyle\pm 1.41}$ $0.1$ $2.88_{\scriptstyle\pm 1.28}$ $2.93_{\scriptstyle\pm 1.23}$ $0.05$
+Usability $3.3_{\scriptstyle\pm 1.43}$ $3.4_{\scriptstyle\pm 1.42}$ $0.1$ $2.75_{\scriptstyle\pm 1.39}$ $2.80_{\scriptstyle\pm 1.22}$ $0.05$
+Table 2: Expert qualitative evaluation of the explanations from Gspell and GNNExplainer on Cora and Amazon-Product.
+$M_{1}$ denotes the mean scores for GNNExplainer and $M_{2}$ the mean scores for Gspell.
+Amazon Cora WikiCS
+F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$ F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$ F (%) $\uparrow$ C (%) $\uparrow$ R (%) A (%) $\downarrow$ S $\downarrow$
+GNNExplainer 94.5 14.7 6.9 - 5.71 96.5 15.8 7.8 - 17.91 99.0 13.1 13.1 - 1930.89
+Gspell 91.1 8.2 4.8 3.0 4.25 86.5 5.8 1.2 7.8 2.07 86.5 4.20 0.0 4.7 12.80
+Gspell+ 92.0 8.3 3.5 5.0 4.24 85.8 6.2 1.5 8.2 2.10 86.3 3.8 0.2 5.8 12.44
+Table 3: Perturbation-based evaluation of faithfulness.
+Gspell shows a superior trade-off between faithfulness and conciseness, achieved by isolating critical nodes for human-centered explanations in natural language.
 Baselines.
-To compare our method’s performance with existing approaches, we consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph.
+We consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph.
 This represents the minimal possible context and serves as a lower bound on explanation size.
@@
 The results are averaged over 5 samples.
-(3-5) Graph Explainers: We compare Logic with GNNExplainer Ying et al.
+(3-5) Graph Explainers: We compare Gspell with GNNExplainer Ying et al.
 (2019a), PGExplainer Luo et al.
-(2020), subgraph-based explainers for node classification tasks, which do not involve any LLM-based component.
-We add the recent LLM-based graph explanation method, TAGExplainer Pan et al.
+(2020) as subgraph-based explainers, which do not involve any LLM-based component.
+We also include the recent LLM-based graph explanation method, TAGExplainer Pan et al.
 (2024).
-To the best of our knowledge, TAGExplainer is the only existing LLM-based explanation approach specifically designed for TAGs.
-5.3 Quantitative Retults
-Table 1 presents the results.
-One major observation is that on all the datasets, the Node baseline (target node as the singleton explanation subgraph) is considerably powerful, denoting how informative the textual attributes of each node are to guide the GNN’s predictions.
-Assessing based on fidelity, GNNExplainer, achieves the best performance over most of the datasets, except for Cora, on which Logic achieves better results.
-However, as discussed before, fidelity alone does not reflect the superiority of a method.
-Essentially, conventional GNNs (GCN in our case) score a sample, i.e., a node, based on the combination of its own information and the aggregated information of its neighbors, meaning that the context surrounding a node highly impacts its prediction score.
-Therefore, a GNN explainer that explains a target node with its whole induced subgraph achieves 100% fidelity.
-Meanwhile, the majority of nodes in the same neighborhood of graphs with a high homophily ratio belong to the same category.
-Thus, selecting a comparably large number of neighbors of a target node on these types of graphs can most probably guarantee the same representation for the target node as if the whole induced subgraph is selected.
-Logic achieves a strong balance between fidelity and explanation size across all four datasets.
-Notably, Logic matches or exceeds the fidelity of traditional explainers like GNNExplainer, outperforming it on Cora, while producing substantially smaller explanation subgraphs, particularly on Liar and Amazon.
-This result is especially significant for TAGs, where textual node attributes carry rich semantic information; Logic effectively distills these semantics into compact, interpretable explanations without compromising on alignment with the original GNN predictions.
-In contrast, baseline methods such as Random or Node either inflate the explanation size or overly restrict the context.
-These results validate our design choice of using LLMs guided by projected GNN embeddings to construct context-aware yet concise explanation subgraphs, offering a superior trade-off between interpretability and predictive fidelity.
-We present additional experiments for different GNN architectures and LLMs in Appendix E.
-Table 2: Expert qualitative evaluation of the explanations from Logic and GNNexplainers on Cora. $M_{1}$ denotes the mean scores for GNNexplainer and $M_{2}$ the mean scores for Logic.
-Dimension $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta(M_{2},M_{1}$)
-Understandability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
-Trustworthiness $2.7_{\scriptstyle\pm 0.8}$ $3.1_{\scriptstyle\pm 0.8}$ $0.4_{\scriptstyle\pm 0.8}$
-Insightfulness $3.3_{\scriptstyle\pm 1.1}$ $3.1_{\scriptstyle\pm 1.1}$ $0.2_{\scriptstyle\pm 1.1}$
-Satisfaction $2.2_{\scriptstyle\pm 0.9}$ $2.7_{\scriptstyle\pm 0.9}$ $0.5_{\scriptstyle\pm 0.9}$
-Confidence $0_{\scriptstyle\pm 0.9} 3.3$$3_{\scriptstyle\pm 0.9}$ $0.3_{\scriptstyle\pm 0.9}$
-Convincingness $3.3_{\scriptstyle\pm 1.2}$ $3.7_{\scriptstyle\pm 1.2}$ $0.4_{\scriptstyle\pm 1.2}$
-Communicability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
-Usability $3.3_{\scriptstyle\pm 0.4}$ $3.5_{\scriptstyle\pm 0.4}$ $0.2_{\scriptstyle\pm 0.4}$
+5.2 Quantitative Results
+Table 1 shows that the Node baseline is relatively strong, highlighting the critical role of textual attributes in GNN predictions.
+While GNNExplainer often reaches the highest fidelity, this metric alone can be misleading.
+Since GNNs aggregate neighborhood data, high fidelity can be trivially achieved by simply outputting the entire induced subgraph.
+In homophilic graphs, where neighbors often share the same category, selecting a broad neighborhood effectively replicates the GNN’s original context.
+This is evidenced by the Random baseline, which achieves fidelity scores comparable to GNNExplainer.
+In these settings, attaining high fidelity is trivial if the explanation subgraph is large.
+Across all datasets, Gspell achieves a superior fidelity-size trade-off: despite marginally lower fidelity than GNNExplainer and the LLM baseline, it produces significantly smaller subgraphs.
+Unlike Random or Node baselines, which either inflate size or restrict context, Gspell leverages projected GNN embeddings to distill rich textual semantics into concise, interpretable explanations.
+Comparing Gspell+ to Gspell highlights the impact of the InfoNCE objective: while it further reduces explanation size, it occasionally lowers fidelity.
+The objective’s effectiveness scales with the number of negative samples; for instance, it outperforms on Amazon-Product compared to Cora because a higher class count provides a tighter mutual information bound.
+Perturbation-based Evaluation.
+To further assess the faithfulness of the generated explanations, we adopt perturbation-based evaluation.
+[…]
+Unlike baselines that provide binary-mask explanations, Gspell ’s 3-way classification—Support, Oppose, and Neutral—uniquely enables the Alignment metric to differentiate between contradicting evidence and irrelevant noise to detect hallucinations.
+Across all datasets, Gspell’s Alignment scores are low, particularly on Amazon, confirming successful detection of irrelevant nodes.
 5.3 Qualitative Evaluation
-Results on Cora.
-We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix F for a full description of the evaluation dimensions.
-An expert reviewed the explanations from GNNexplainer and from our method over 10 node classification explanation tasks, and the evaluation framework is developed in Appendix G.
-The results are presented in Table 5.3.
-We observe that our method appeared more trustworthy and convincing to our expert, which may be explained by the semantic analysis abilities of the LLM of our graph explaining method.
-Concurrently, the explanations were also less insightful, as this may have limited the LLM in its exploration of topics and application not directly limited to the subject of the research paper.
-Visualization on Amazon dataset.
-Figure 2 illustrates a representative example of Logic output on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation.
-In this instance, the target (bold) node corresponds to a product (ID 7), predicted to belong to the “Sports & Outdoors” category.
-Logic analyzes the local neighborhood of the target node and translates the latent GNN-based representations into natural language explanations by leveraging the LLM’s language generation capabilities, interleaved with soft GNN prompts.
-Logic produces node-wise rationales from the computation tree of the target node, explicitly stating for each neighbor a brief semantic summary along with a binary support judgment (YES/NO) regarding its influence on the target’s classification.
-Table 3: Performance comparison for the ablation study.
-LLM denotes prompting the LLM with words only and no projected embeddings.
-LLM+Pr is LLM with a projector.
-LLM with projector and post-processing (denoted by LLM+Pr+Po) is our method.
-Cora WikiCS Liar Amazon
-Fidelity Size Fidelity Size Fidelity Size Fidelity Size
-LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.2
-LLM+Pr 92.2 % 16.0 80.1% 3.7 95.0% 10.1 86.5% 1.7
-LLM+Pr+Po 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 4.9
-5.5 Ablation Study
-To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing.
-The results are in Table 3.
-We observe that both fidelity and explanation size drop across all datasets when we replace the words in the prompt with the projected GNN embeddings.
-This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning.
-We also observe that adding the post-processing step increases fidelity, at a reduced cost in size compared to the method without a projector, on the WikiCS dataset.
-From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
+Expert Assessment.
+We asked experts to evaluate the explanations generated by GNNExplainer and Gspell on the Amazon-Product and Cora datasets over 10 node classification tasks.
+We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix E for a full description of the evaluation dimensions.
+The evaluation framework and setup are presented in Appendix F.
+$k=1$ $k=5$ $k=10$ $k=20$ $k=50$
+Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
+Amazon Gspell 92.0 4.22 91.1 4.25 88.0 1.71 85.0 1.17 82.5 1.00
+Gspell+ 90.5 4.45 92.0 4.24 86.0 1.45 85.0 1.13 81.0 1.01
+Liar Gspell 100.0 1.13 100.0 1.11 100.0 1.11 100.0 1.00 OOM OOM
+Gspell+ 100.0 1.14 100.0 1.10 98.4 1.10 100.0 1.00 OOM OOM
+Cora Gspell 86.2 1.98 86.5 2.07 87.8 2.04 85.4 1.81 84.4 1.33
+Gspell+ 85.2 1.87 85.8 2.10 88.0 2.17 87.4 2.17 84.8 1.30
+Table 4: Effect of the number of tokens ($k$) on the performance of Gspell and Gspell+.
+As Table 2 shows, Gspell’s explanations on Cora were more trustworthy and convincing, likely due to the LLM component’s semantic understanding.
+However, these explanations were seen as less insightful, suggesting more limited exploration beyond the article’s main subject.
+On Amazon-Product, Gspell provided more insightful explanations than GNNExplainer by capturing richer semantic links across product categories, but it was rated lower in trustworthiness and convincingness, indicating less consistency and reliability.
+Overall, our method performs better on Cora, where its explanations were viewed as more trustworthy, convincing, and communicable.
+On the more challenging Amazon-Product dataset, GNNExplainer retained an edge in clarity and user confidence, suggesting that dataset complexity and semantic richness shape the relative strengths of LLM-based explanations.
+Real Example.
+Figure 2 illustrates a prompting example of Gspell on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation.
+In this instance, the target node corresponds to a product (ID 160), predicted to belong to the “Sports & Outdoors” category.
+Gspell leverages latent GNN representations via soft prompts to benefit from the LLM’s language-generation capabilities while accounting for the local neighborhood of the target node.
+Gspell generates node-level rationales from the target node’s computation tree, providing for each neighbor a binary support judgment (YES/NO) indicating whether it supports the target’s classification.
+Also, it provides a general summary that highlights common patterns across the target’s context, helping to explain the model’s prediction.
+6 Ablation Study and Runtime Analysis
+6.1 Effect of the projection tokens
+Without accounting for GNN internals (embeddings), asking an LLM to generate explanations only based on the input and the predictions of the GNN is theoretically wrong.
+The main reason is that one can directly substitute the predictive GNN with any arbitrary function that makes predictions based on different reasoning (logic) than the GNN, while still making the same predictions.
+Therefore, we use the soft prompt generated by the projector for the faithfulness of the explanations.
+Soft prompts channel information from the GNN to the LLM, with the token count, $k$, being a critical factor here.
+Insufficient tokens constrain the information channel, causing the LLM to ignore GNN-derived context in favor of textual priors, thereby increasing hallucination rates.
+Conversely, a large number of tokens can disrupt coherence during generation.
+If the soft prompt distribution aligns poorly with the LLM’s embedding space—often due to a suboptimal projection function–increasing $k$ injects out-of-distribution noise into the model.
+This results in degraded generation quality and increases computational overhead.
+We perform an ablation study over different token counts ($k$), shown in Table 4.
+We observe that $k$ acts as a sparsity regulator.
+On the complex Amazon dataset, low $k$ yields high fidelity with larger explanations, while high $k$ significantly reduces size at the cost of fidelity.
+This supports our hypothesis that excessive soft tokens inject noise into the LLM.
+We find an optimal balance at $k=5$, achieving high fidelity with reduced explanation size.
+In contrast, Liar, a more heterogeneous dataset, benefits from a higher $k$ as it likely requires more tokens to encode complex relationships than standard homogeneous graphs.
+Conversely, Cora exhibits high robustness in both fidelity and sparsity across all $k$ values, likely due to its higher homophily.
+6.2 Runtime
+We provide the details of running times in Table 5, comparing our method with all baselines.
+Although Gspell is more time-consuming than other baselines, we note that this is justified by the fact that Gspell, unlike the fast baselines, e.g., PGExplainer and TAGExplainer, generates descriptive, informative explanations in natural language that are understandable by humans.
+Amazon Cora Liar WikiCS
+GNNExplainer 1.45 1.69 1.79 3.76
+PGExplainer 0.02 0.03 0.79 0.80
+TAGExplainer 0.05 0.13 0.52 1.57
+LLM 33.74 18.21 12.80 61.67
+Gspell (Ours) 39.65 30.77 42.41 76.76
+Table 5: Inference runtime per node for explanation generation measured in seconds.
 7 Conclusions
-In this work, we propose Logic, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs.
-By projecting GNN embeddings into the LLM space, Logic enables context-aware, natural language rationales without relying on traditional GNN explainer modules that often need training.
-Our approach produces concise explanation subgraphs with high fidelity and semantic clarity, outperforming prior methods.
-Logic demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
+In this work, we propose Gspell, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs.
+By projecting GNN embeddings into the LLM space, Gspell enables context-aware, natural language rationales without relying on an external GNN explainer.
+Our approach achieves a superior trade-off between conciseness and fidelity, as well as greater semantic clarity, compared to prior methods.
+Gspell demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
+Future Work.
+Although Gspell+ incorporates an InfoNCE loss to enhance faithfulness, it does not consistently outperform Gspell.
+Moreover, the other projector objectives focus on aligning GNN and LLM embedding spaces rather than directly optimizing explanation quality.
+Designing more principled, explanation-centered objectives is an important direction for future work.
+Additionally, as the computation tree grows, the LLM struggles to capture the full graph context.
+While batch-wise prompting mitigates this issue, it limits global reasoning.
+Developing prompt compression techniques that preserve structural information could improve the quality of the generated explanations.
+Acknowledgment
+The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Texas Advanced Computing Center (TACC) Vista for contributing to this research.
 8 Ethical considerations
 In this work, we have built methods for generating explanations for graph neural network predictions. We do not foresee any ethical issues from our study.
+Potential Risks.
+While Gspell aims to improve the interpretability of GNN predictions, the generated explanations rely on LLM outputs and may not always faithfully reflect the underlying model’s true decision process.
+This could lead to over-trust in the explanations.
+Also, the use of LLMs introduces the risk of hallucinated or misleading reasoning, despite our mitigation strategies.
+We encourage careful human oversight when deploying such explanation systems in practice.
 9 Limitations
 We present some limitations of our work:
-  • Our method projects GNN embeddings into the LLM embedding space and interleaves them with text to form hybrid prompts.
-However, batching multiple nodes leads to degraded performance and hallucinations, likely due to interference between soft prompts.
-  • While Logic achieves high empirical fidelity, it does not provide formal guarantees of faithfulness to the GNN’s internal decision logic, as the explanation relies on the LLM’s interpretation of projected embeddings rather than directly tracing GNN computation paths.
-  • Although our method incorporated a projector to align the embeddings of the GNN with the LLM space of token embeddings, the objectives we introduce for training the projector are not directly tailored for the task of explanation.
-More optimal objectives that adhere to the explanation task could bring substantial benefits.
+  • Our method combines projected GNN embeddings with text in hybrid prompts.
+When the computation tree is large, processing the full context can lead to hallucinations, as the LLM struggles to distinguish among multiple soft prompt representations, while batch-wise prompting limits its ability to capture a global view.
+  • The explanations generated by Gspell are produced by an LLM conditioned on projected embeddings rather than explicitly tracing the GNN’s computation.
+Consequently, they may not reflect the true causal factors underlying the model’s prediction, so we do not provide formal guarantees of causal faithfulness.
+  • Although the projector aligns GNN embeddings with the LLM embedding space, its training objectives are not directly tailored for explanation.
+While we explore an InfoNCE-based objective to improve faithfulness, it does not consistently yield gains, suggesting that more effective explanation-specific objectives are needed.
 Appendix A Additional Related Work
 Here we provide additional related works that are on the integrated methods of LLMs and GNNs.
-The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017b); Hamilton et al.
+The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017a); Hamilton et al.
 (2017).
@@
 LLMs as prefix and postfix.
-First, there are methods that utilize LLMs at the prefix modules of their frameworks for task and feature representation Liu et al.
+First, there are methods where LLMs are used in the prefix modules for task and feature representation Liu et al.
 (2024a); Fang et al.
@@
 (2024); Tang et al.
-(2024), which is better suited for capturing a graph’s topology and its intricate relationships and uses LLMs for prediction.
-This approach also offers flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al.
+(2024), which is better suited for capturing the graph topology and its intricate relationships and uses LLMs for prediction.
+This approach also has some flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al.
 (2023); Chen et al.
@@
 (2024).
-Approaches following this paradigm benefit from the generalization capabilities of LLMs at both input and output levels, and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts.
+Approaches following this paradigm benefit from LLMs’ generalization capabilities at both the input and output levels and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts.
 Finally, there is a group of works which employ Gnns and LLMs jointly through soft prompting, which necessitates alignments between their embeddings while enhancing LLMs reasoning by the Gnn captured graph structural features Liu et al.
@@
 (2023).
+However, these works do not address the GNN explanation problem.
 Explainers for GNNs.
@@
 For more details, please refer to this survey on GNN explainers Kakkad et al.
-(2023b).
-Table 4: Statistics of the graph datasets.
+(2023).
 Dataset # Nodes # Edges # Categories # Features
@@
 Liar 13293 32443 7 387
+Table 6: Statistics of the graph datasets.
 Appendix B Datasets
+Dataset Description & Relevance Why It Fits Our Work
+CORA Citation network of CS papers with bag-of-words features. Standard benchmark for node classification on text-rich graphs.
+WIKICS Wikipedia CS articles linked by hyperlinks, with GloVe embeddings. Large, semantic-rich graph ideal for explanation evaluation.
+LIAR Fake news detection graph combining statements, speakers, topics. Challenging, heterogeneous graph testing method adaptability.
+AMAZON Product co-purchase network with 47 categories. Real-world, diverse e-commerce graph to test scalability.
+Table 7: Dataset Descriptions and Relevance to Our Work
 B.1 Datasets Details
@@
 The nodes are classified into one of seven categories.
-Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors.
+Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing-based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors.
 Each node’s feature is a 2000-dimensional bag-of-words representation of the keywords in the paper that it corresponds to.
@@
     WikiCS: It is a text-attributed graph dataset, derived from the Wikipedia platform, widely used for node classification tasks.
-The nodes correspond to Wikipedia page descriptions of different computer science topics, and edges between nodes represent hyperlinks from one article to another.
+The nodes correspond to Wikipedia page descriptions of different computer science topics, and the edges represent hyperlinks from one article to another.
 Each node in the dataset belongs to one of 10 categories.
-The node classification task is to correctly predict a node’s label, and the edges are considered to be undirected.
-Each node’s feature vector is a GloVe embedding of size 300, computed based on the text associated with the nodes.
+The node classification task is to correctly predict a node’s label, and edges are undirected.
+Each node’s feature vector is a 300-dimensional GloVe embedding computed from the text associated with the node.
   3. 3.
-    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about.To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set.
+    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about.
+To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set.
 Each node is embedded with a 384-dimensional BERT representation and augmented with a 3-dimensional one-hot vector indicating its type.
@@
 Statement nodes retain their ground-truth labels for fake news classification (ranging from pants-fire to true), while speaker and topic nodes are assigned a dummy label.
-We also preserve a node-level string attribute—containing the raw statement, speaker name, or topic—for use in explanation and visualization.
+We also preserve a node-level string attribute containing the raw statement, speaker name, or topic for use in explanations and visualizations.
   4. 4.
-    Amazon-Product: It is a network with nodes representing different kinds of products on Amazon and edges connecting co-purchased products.
-The nodes belong to one of 47 different categories of products.
+    Amazon-Product: It is a network with nodes representing different product categories on Amazon and edges connecting co-purchased products.
+The nodes belong to one of 47 product categories.
 We use only a subset of the Amazon-Product   dataset, consisting of the first 1000 products and their co-purchase edges, as the entire dataset is very large, and call it Amazon   in our work.
-The node classification task is to predict a product’s category based on its own as well as its neighbours’ features.
+The node classification task is to predict a product’s category based on its own features and those of its neighbors.
+For nodes, bag-of-words representations are derived from node textual attributes.
 B.2 Relevance of datasets
-Table 5: Dataset Descriptions and Relevance to Our Work
-Dataset Description & Relevance Why It Fits Our Work
-CORA Citation network of CS papers with bag-of-words features. Standard benchmark for node classification on text-rich graphs.
-WIKICS Wikipedia CS articles linked by hyperlinks, with GloVe embeddings. Large, semantic-rich graph ideal for explanation evaluation.
-LIAR Fake news detection graph combining statements, speakers, topics. Challenging, heterogeneous graph testing method adaptability.
-AMAZON Product co-purchase network with 47 categories. Real-world, diverse e-commerce graph to test scalability.
-The datasets used in our work are selected based on their relevance to graph explainability tasks, allowing us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features.
-The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake-news detection, and e-commerce, which reflect both real-world complexity and common challenges faced in explainability.
+The datasets used in our work are selected for their relevance to graph explainability tasks, enabling us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features.
+The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake news detection, and e-commerce, and reflect both real-world complexity and common challenges in explainability.
 The characteristics of each dataset make them suitable for evaluating the performance of our explanation method.
 Table 7 summarizes the key features of the datasets and explains their relevance.
-Appendix C Additional Implementation Details
-We discuss additional details related to the training of the base GNN.
-C.1 GNN training
-For all datasets, we implement a GCN with 3 convolutional layers, with the input and output channel dimensions chosen according to each dataset’s feature dimension and number of classes, respectively.
+Appendix C Extended Experiments
+C.1 GNN Architecture
+We extend Table 1 of the paper with additional experiments for Amazon using two other standard GNN architectures: GIN Xu et al. and GAT Veličković et al. (2017). The results are reported in Table 8.
+Method Model Fidelity (%) Size
+GNNExplainer GAT 100.0 10.57
+PGExplainer 71.8 4.45
+TAGExplainer 63.1 1.03
+Random 89.9 7.28
+LLM 96.5 4.92
+Gspell 86.8 3.79
+GNNExplainer GIN 100.0 10.57
+PGExplainer 87.2 5.15
+TAGExplainer 30.9 1.03
+Random 84.7 7.28
+LLM 86.0 4.90
+Gspell 73.7 3.06
+Table 8: Performance comparison on the Products dataset using different GAT and GIN as two other architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
+With GCN architecture, we observed that our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance.
+However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (73.7%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation.
+In contrast, GNNExplainer maintains the highest fidelity across all architectures, but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes.
+The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise.
+These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structure-based models.
+Llama 3.1 8B Instruct Mistral 7B Instruct v0.2 GPT-Neo 2.7B Phi-3 Mini 4k Instruct Pythia 2.8B
+Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
+LLM 86.0 4.90 88.0 5.14 84.0 2.33 88.0 5.17 84.0 2.34
+Gspell 91.1 4.25 83.0 1.63 80.0 1.32 80.0 1.47 80.0 1.12
+Table 9: Results on Products dataset using different pretrained LLMs (for pretrained GCN). Higher fidelity and lower size are better.
+C.2 LLM Backbone
+We also evaluate our method across different LLM models and present the results in Table 9.
+The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger.
+In contrast, our method consistently generates more compact explanations—typically around size 1 —while maintaining comparable fidelity.
+Notably, Llama 3.1 and Mistral backbones show strong fidelity across both methods, whereas smaller models such as GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall.
+These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
+C.3 Framework Component Analysis
+To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing.
+The results are in Table 10.
+We observe that both fidelity and explanation size decrease across all datasets when we replace the prompt words with the projected GNN embeddings.
+This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning.
+We also observe that adding the post-processing step increases fidelity at a lower cost in size than the method without a projector on the WikiCS dataset.
+From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
+Amazon Cora Liar WikiCS
+Fidelity Size Fidelity Size Fidelity Size Fidelity Size
+LLM 86.0 4.90 94.5 3.27 100.0 2.22 91.4 21.14
+LLM+Pr 85.8 3.2 94.5 3.27 100.0 1.10 80.2 3.7
+LLM+Pr+Po 91.1 4.25 86.5 2.07 100.0 1.11 86.5 12.8
+Table 10: Performance comparison for the ablation study.
+LLM denotes prompting the LLM with words only and no projected embeddings.
+LLM+Pr is LLM with a projector.
+LLM with projector and post-processing (denoted by LLM+Pr+Po) is our method.
+Appendix D Additional Implementation Details
+We provide additional details of the experiments as follows:
+D.1 Experimental Setup
+Datasets.
+For all datasets, we use the graph topology as provided, without any modification or filtering of nodes or classes.
+For AMAZON-PRODUCT, we first randomly subsample 1000 nodes from the full dataset.
+For all datasets, we use a consistent split protocol with 60% of nodes for training, 10% for validation, and 200 held-out nodes for testing.
+Across all datasets, we restrict computation trees to at most 2 hops from each node.
+Models.
+We use a GCN Kipf and Welling (2017b) as the base GNN model for all datasets.
+The GCN consists of 3 convolutional layers, with the final layer used for classification.
+The model is trained for 100 to 400 epochs (tune based on validation set for each dataset) using the Adam optimizer with a batch size of 64.
 For the hidden dimension, we choose either $64$ or $512$, depending on the nodes’ feature dimension.
-The model is trained with the Adam optimizer Kingma and Ba (2014) and a learning rate of 0.001 for 400 epochs.
-C.2 GNN accuracy
-We give the GNNs accuracy for the different folds in Table 6.
-Table 6: Accuracy of GCN on the node classification task for the Cora, Wiki CS, Liar and Amazon products datasets.
+We provide the GNN’s accuracy where GCN is used as the model architecture for the different folds in Table 11.
 Dataset Training Validation Testing
@@
 Liar 0.42 0.30 0.29
-Appendix D Prompt and generated responses examples
-In this section we present a detailed prompt construction in D.1 and a complete example of such built prompt with its input in D.2.
-D.1 Example of Full Prompt Construction for Logic
-We have provided an example of the full prompt used in Logic below.
-The boldface texts denote where the embeddings are injected into the prompt.
-An example response is also illustrated in Figure 3 in the paper.
-Prompt:
-You are an expert research assistant.
-Your goal is to assess whether the classification of a target product is supported by its co-purchase neighborhood by analyzing Amazon product reviews and its co-purchase neighborhood to understand why it has been classified under a specific category.
-The categories to distinguish from are Home & Kitchen, Health & Personal Care, Beauty, etc.
-Target Product ID: 0
-Predicted Category: Clothing, Shoes & Jewelry
-Target Product Embedding Representation:
-[…]
-Neighboring Products in the Co-purchase Network:
+Table 11: Accuracy of GCN for the node classification task on four datasets.
+For main experiments, we use Meta’s Llama-3.1-8B-Instruct model from the HuggingFace library as a frozen LLM.
+Projector.
+The projector is implemented as a two-layer MLP with ReLU activation.
+The hidden dimension is set to 4x the GNN embedding dimension.
+The output layer maps the GNN embedding to a vector of size $k\cdot h$, where $k$ is the number of soft prompt tokens and $h$ is the LLM embedding dimension.
+The mean-pooled representation is used for alignment objectives.
+We fix $k=5$ across all datasets.
+The projector is trained separately for each dataset using full-batch gradient descent.
+We use the Adam optimizer, with the learning rate selected via random search over $\{0.001,0.0005\}$ based on validation performance.
+The number of training epochs is selected from $\{100,200,400\}$ using validation.
+The loss weights are selected via random search over $\{0.1,0.5,0.9\}$ on the validation set.
+The contrastive temperature is fixed to $\tau=0.2$.
+Explanation construction.
+For each node, we construct explanation subgraphs by including all supporting nodes and excluding all opposing nodes.
+LLM inference.
+We use the default HuggingFace generation settings for the LLM, without modifying temperature or sampling parameters, and set the maximum number of generated tokens to 1024.
+Prompts are formatted using the chat template when supported by the model, and a fixed prompt template is used across all datasets.
+Baselines and evaluation protocol.
+We compare Gspell against standard GNN explainers including GNNExplainer, PGExplainer, and TAGExplainer.
+We use official implementations and PyTorch Geometric implementations where applicable, without additional hyperparameter tuning.
+All methods are evaluated under the same GNN backbone for fairness. Experiments are repeated over 5 random seeds, and we report the mean and standard deviation of all metrics.
+We discuss additional details related to the training of the base GNN.
+D.2 Additional example on Amazon-Product
+A key advantage of Gspell is its ability to increase human understandability and trustworthiness by providing explanation of explanations.
+We give an additional example on Amazon-Product in Figure 3.
+The “Reasoning” section of the LLM response is where the LLM provides a descriptive reasoning on how the supporting neighbors can explain the prediction of the target node.
+On the other hand, the previous GNN explainers simply generate a subgraph that is more abstract and harder to understand without some level of background knowledge.
+LLM Prompt You are analyzing Amazon product reviews and their co-purchase neighborhoods to understand why it has been classified under a specific category.
+Your task is to determine which of the neighboring products does (not) support the classification of the target product.
+Target Product ID: 160 Predicted Category: Health & Personal Care
+Product Embedding Representation:
+<<BEGIN TARGET KEYWORDS>> PROJECTED EMBEDDING <<END TARGET KEYWORDS>>
+Neighboring Products:
 Each Product below is described by keywords.
-  • Product 129:
-[…]
-  • Product 130:
-[…]
-Instructions:
-You are given a target product and a list of neighboring products, each described by keywords.
-For each neighboring product:
-  1. 1.
-    Write one sentence summarizing the main topics or ideas captured in its keywords.
-  2. 2.
-    Clearly state whether this product supports the classification of the target product into category ‘Clothing, Shoes & Jewelry’.
-Use the following format for each neighbor:
-Product {ID}:
-Summary: One sentence summary of the product’s keywords.
-Support: YES or NO — Does this product support the classification into ‘Clothing, Shoes & Jewelry’?
-Base your reasoning only on the keywords and proximity to the target product.
-Start your analysis below:
-D.2 Sample of the prompt and output for the Amazon-Product dataset
-We provide a sample of the prompt and a sum up of the output we obtain (without the supporting sentence for each product) in Figure 3.
-Figure 3: An illustration of the prompting template and LLM explanation of Logic for the Amazon dataset.
-The left block shows the input prompt to the LLM for a sample target node (target node id 3).
-The right block denotes the LLM response explaining the prediction of the GNN for the target node.
-Appendix E Additional Experiments
-We extend Table 1 of the paper with additional experiments for Amazon with TAGExplainer Pan et al.
-(2024), PGExplainer Luo et al.
-(2020) and with the GIN Xu et al.
-and GAT Veličković et al.
-(2017) architectures.
-The results are reported in Table 7.
-Table 7: Performance comparison on the Products dataset using different GNN architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
-GCN GAT GIN
-Fidelity Size Fidelity Size Fidelity Size
-Node 73.8% 1.00 73.2% 1.00 39.6% 1.00
-Random 84.4% 7.24 84.2% 7.28 90.2% 7.25
-GNNExplainer 95.5% 5.63 94.5% 5.615 88.0% 5.65
-PGExplainer 87.2% 5.46 88.3% 5.76 87.2% 5.46
-TAGExplainer 73.8% 1.37 71.8% 1.43 42.3% 1.27
-LLM 86.5% 5.07 91.0% 5.00 86.5% 5.18
-Our method 87.0% 1.3 81.5% 1.45 60.0% 1.52
-Our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance for the GCN with 87.0% fidelity at just 1.3 in size.
-However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (60.0%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation.
-In contrast, GNNExplainer maintains the highest fidelity across all architectures but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes.
-The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise.
-These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structural-based models.
-Table 8: Results on Products dataset using different pretrained LLMs (for pretrained GCN). Higher fidelity and lower size are better.
-Llama 3.1 8B Instruct Mistral 7B Instruct v0.2 GPT-Neo 2.7B Phi-3 Mini 4k Instruct Pythia 2.8B
-Fidelity Size Fidelity Size Fidelity Size Fidelity Size Fidelity Size
-LLM 86.5% 5.07 88.0% 5.14 84.0% 2.33 88.0% 5.17 84.0% 2.34
-Our method 87.0% 1.30 83.0% 1.63 80.0% 1.00 80.0% 1.00 80.0% 1.00
-The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger.
-In contrast, our method consistently generates more compact explanations—typically around size 1, with a moderate decrease in fidelity of roughly 4–7 percentage points depending on the model.
-Notably, Llama 3.1 and Mistral backbones show strong fidelity for both methods, whereas smaller models like GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall.
-These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
-Appendix F Definition of the qualitative metrics
+Product 200: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 201: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 202: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 203: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 204: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 205: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 206: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 207: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 208: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 209: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+Product 210: <<BEGIN KEYWORDS>> PROJECTED EMBEDDING <<END KEYWORDS>>
+LLM Response P200 Summary: Home care and hygiene products.
+NO P201 Summary: Personal grooming and beauty tools.
+NO P202 Summary: Health and wellness supplements.
+YES P203 Summary: Fitness and exercise equipment.
+NO P204 Summary: Skincare and beauty products.
+YES P205 Summary: Health and wellness journals.
+YES P206 Summary: Personal care and hygiene for specific needs.
+YES P207 Summary: Health and wellness supplements.
+YES P208 Summary: Fitness and exercise accessories.
+NO P209 Summary: Health and wellness products for specific needs.
+YES P210 Summary: Health and wellness journals.
+YES Reasoning: The common pattern among the supporting Products is that they are all related to health and wellness, skincare, and personal care, which are all categories within the Health & Personal Care category.
+Instructions For each neighboring product: 1.
+Summarize the similarity between supporting products in one sentence.
+2.
+Clearly state whether this product supports the classification of the Target Product into category Health & Personal Care.
+Format:
+Product <ID>:
+Summary: <One sentence>.
+Support: YES or NO — Does this product support classification into Health & Personal Care?
+Base reasoning only on the keywords and proximity to the target product.
+Figure 3: Left: prompt with category and embeddings highlighted. Right: model response with summaries, YES/NO verdicts, and reasoning. Below: instructions shown separately.
+Appendix E Definition of the qualitative metrics
 We adapt the different dimensions of the framework in Cedro and Martens (2025) to the context of Cora to clarify their definitions, and provide the definition as given to the expert.
@@
     Usability. Explanations should support practical tasks such as interpreting predictions, or improving model performance.
-Appendix G Experimental Setup for the human evaluation of M1 and M2 Scores
+Appendix F Experimental Setup for the human evaluation of M1 and M2 Scores
+The human evaluation was conducted by the authors, who served as domain experts.
+For the qualitative experiments, one expert reviewed the Cora dataset, and four experts reviewed the Amazon-Product data, evaluating the explanations generated by GNNexplainer and Gspell over 10 node classification explanation tasks.
 We clarify the experimental setup used to obtain the M1 and M2 scores reported in Table 2.
 A) Data Preparation
-We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained. For each article, we generated explanations using both LOGIC and GNNExplainer. Each expert was provided with:
+We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained.
+For each article, we generated explanations using both Gspell and GNNExplainer.
+To mitigate potential bias, explanations from these methods were randomized and presented in a blind manner, without revealing their source.
+Each expert was provided with:
   • The original article text and the GNN’s predicted label.
@@
 B) Evaluation Procedure
-  • Before the evaluation, we provided experts with definitions of the evaluation metrics (as described in Appendix F) and asked them to consider how to apply each metric to content from the Cora dataset.
+  • Before the evaluation, we provided 3 experts with the evaluation metrics’ definitions and asked them to consider how to apply each metric to content from the Cora dataset.
   • During the evaluation, each expert received the materials described above and independently rated each explanation on a scale from 1 to 5 for each metric.
 C) Compilation of Results
-The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Logic. These results are reported in Table 5.3.
+The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Gspell. These results are reported in Table 2.